### v3

# TabRep: a Simple and Effective Continuous Representation for Training Tabular Diffusion Models

Abstract
Diffusion models have been the predominant generative model for tabular data generation. However, they face the conundrum of modeling under a separate versus a unified data representation. The former encounters the challenge of jointly modeling all multi-modal distributions of tabular data in one model. While the latter alleviates this by learning a single representation for all features, it currently leverages sparse suboptimal encoding heuristics and necessitates additional computation costs. In this work, we address the latter by presenting TabRep, a tabular diffusion architecture trained with a unified continuous representation. To motivate the design of our representation, we provide geometric insights into how the data manifold affects diffusion models. The key attributes of our representation are composed of its density, flexibility to provide ample separability for nominal features, and ability to preserve intrinsic relationships. Ultimately, TabRep provides a simple yet effective approach for training tabular diffusion models under a continuous data manifold. Our results showcase that TabRep achieves superior performance across a broad suite of evaluations. It is the first to synthesize tabular data that exceeds the downstream quality of the original datasets while preserving privacy and remaining computationally efficient. Code is available at: https://github.com/jacobyhsi/TabRep.
1 Introduction
Tabular data are ubiquitous in data ecosystems of many sectors such as healthcare, finance, and insurance [11, 35, 12]. These industries utilize tabular data generation for many practical purposes, including data augmentation [22], privacy-preserving machine learning [57], and handling sparse, imbalanced datasets [38, 44]. Unlike homogeneous data modalities such as images or text, one notable characteristic inherent to tabular data is feature heterogeneity [32]. Tabular data often contain mixed feature types, ranging from (dense) continuous features to (sparse) categorical features.
Recently, the best-performing tabular generative models are based on diffusion models [19, 50]. For instance, [30, 26, 47] leverage continuous [19, 50] and discrete diffusion [21, 46] to generate tabular data. However, these methods are designed to jointly optimize a multimodal continuous-discrete data representation that complicates the training process. On the other hand, while solutions that optimize a unified continuous representation circumvent this by learning a single representation for all features, they rely on encoding heuristics such as one-hot [24] or learned latent embeddings using a $\beta$-VAE [58]. Unfortunately, one-hot encoding for categorical variables leads to sparse representations in high dimensions, where generative models are susceptible to underfitting [27, 41]. Furthermore, using a latent embedding space [58] requires training the additional VAE embedding model [18, 25] that requires extra computation while being dependent on the quality of the latent space.
Since diffusion-based generative models rely on continuous transformations of denoising score-matching, or invertible mappings between data and latent spaces, designing an effective data representation is key for high-performing diffusion models [5]. In this work, we propose TabRep, a simple and effective continuous representation for training tabular diffusion models. We introduce geometric insights to understand the properties of a continuous data manifold for tabular diffusion models. Then, we craft an abstract and useful representation, that enables diffusion models like DDPM [20] and Flow Matching [31] to capture the posterior distribution efficaciously. TabRep’s representation maintains a dense two-dimensional representation, circumventing the curse of dimensionality. It is also well-suited to model the cardinality of nominal features, ensuring ample separability to distinguish between different categories. Lastly, coupling the cyclical nature of our representation with the separability it offers enables it to facilitate ordinal features. These attributes present desirable characteristics that make it easier for diffusion models to extract meaningful information from a unified continuous tabular data representation. We conduct comprehensive experiments against tabular diffusion baselines across various datasets and benchmarks. The results showcase that TabRep consistently outperforms these baselines in quality, privacy preservation, and computation cost, indicating our superior capabilities to generate tabular data. Our architecture is in Figure 1.
Figure 1: The TabRep Architecture. TabRep transforms and unifies the data space under a continuous regime via the our representation. A diffusion or flow matching process is trained to optimize the denoising network. Once training is completed, samples can be generated through a reverse denoising process before inverse transforming back into their original data representation.
2 Related Work
The latest tabular diffusion models have made considerable progress compared to previous generative models such as VAEs [56] and GANs [56]. This included STaSy [24], which employed a score-matching diffusion model paired with self-paced learning and fine-tuning to stabilize the training process, and CoDi [30], which used separate diffusion schemes for categorical and numerical data along with interconditioning and contrastive learning to improve synergy among features. TabDDPM [26] presented a similar diffusion scheme compared to CoDi and showed that the simple concatenation of categorical and numerical data before and after denoising led to improvements in performance. TabSYN [58] is a latent diffusion model that transformed features into a unified embedding via a $\beta$-VAE [25] before applying EDM diffusion [23] to generate synthetic data. CDTD [36] combines score matching and score interpolation to enforce a unified continuous noise distribution for both continuous and categorical features. TabDiff, couples DDPM [19] and Discrete Masked Diffusion [46, 43] to synthesize tabular data.
3 Tabular Diffusion Models
Unlike images that contain only a single data type, the heterogeneous nature of tabular data requires an intricate mechanism to model continuous and discrete features. We begin with the following preambles where we denote a tabular dataset as $\mathcal{D}=\{\mathbf{z}^{(i)}\}_{i=1}^{N}=\{[\mathbf{x}^{(i)},\mathbf{c}^{(i)}]\}_{i=1}^{N}$, where $N$ represents the number of samples (rows). Each sample consists of continuous (numerical) features $\mathbf{x}^{(i)}\in\mathbb{R}^{D_{\text{cont}}}$ and discrete (categorical) features $\mathbf{c}^{(i)}\in\prod_{j\in\left\{1,\ldots,D_{\text{cat}}\right\}}\{1,\ldots,K_{j}\}$, where $D_{\text{cont}}$ is the number of continuous features, $D_{\text{cat}}$ is the number of categorical features, and $K_{j}$ is the number of unique categories for the $j$-th categorical feature.
3.1 Tabular Diffusion Framework
Traditional Diffusion Models [19, 50, 49] operate under a Markovian noising and denoising process where the forward noising iteratively corrupts the data distribution into a Gaussian distribution via Gaussian convolution, and the reverse denoising learns to transform Gaussian noise back into meaningful data samples. Due to the multi-modality of tabular data, tabular diffusion models have been designed to handle each tabular column as its innate continuous and discrete data type. The joint continuous-discrete forward diffusion process is defined as:
$q(\mathbf{z}_{t}|\mathbf{z}_{0})$ $=q(\mathbf{x}_{t}|\mathbf{x}_{0})\cdot q(\mathbf{c}_{t}|\mathbf{c}_{0}).$ (1)
The reverse joint diffusion process parameterized by $\theta$ is learned and given by:
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c}_{t})\cdot p_{\theta}(\mathbf{c}_{t-1}|\mathbf{x}_{t},\mathbf{c}_{t}).$ (2)
At a high level, the latest state-of-the-art tabular diffusion models [47, 26, 30] are trained on a separate continuous-discrete data representation via continuous Gaussian diffusion [19, 23] and discrete diffusion [46, 21].
Gaussian Diffusion of Continuous Features. The diffusion process of continuous features can be formulated using discrete-time Gaussian diffusion [19, 37]; the noising process is described as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t}|\sqrt{\alpha_{t}^{\mathbf{x}}}\mathbf{x}_{t-1},(1-\alpha_{t}^{\mathbf{x}})\mathbf{I}),$ (3)
where $\alpha_{t}^{\mathbf{x}}$ is a hyperparameter to control the magnitude of the noising process. Per Tweedie’s Lemma [14, 42], the mean of the denoising distribution can be obtained by the score of the distribution:
$\mathbb{E}_{q}[\mathbf{x}_{t-1}|\mathbf{x}_{t}]\!=\!\frac{1}{\sqrt{\alpha_{t}^{\mathbf{x}}}}\left(\mathbf{x}_{t}\!-\!\frac{1-\alpha_{t}^{\mathbf{x}}}{\sqrt{1-\bar{\alpha}_{t}^{\mathbf{x}}}}\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t})\right),$ (4)
in which we can use the approximated score function $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x})\approx\nabla_{\mathbf{x}}\log q(\mathbf{x})$ learned using denoising score matching [50, 53]. This yields the denoising loss function for continuous features:
$\mathcal{L}_{\mathbf{x}}\!=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}_{t}}\!\left[\left\|\nabla_{\!\mathbf{x}_{t}}\!\log q(\mathbf{x}_{t}\!\!\mid\!\!\mathbf{x}_{0})\!-\!\nabla_{\!\mathbf{x}_{t}}\!\log p_{\theta}(\mathbf{x}_{t},\mathbf{c}_{t})\right\|^{2}\right].$ (5)
Discrete Diffusion of Categorical Features. Generating categorical features can be achieved using discrete-time diffusion processes tailored to handle discrete data [21, 7]. The noising process for categorical variables is defined by transitioning between categorical states over discrete timesteps. The reverse process aims to learn the transition probabilities.
Unified Continuous Diffusion. Training diffusion models on a separate continuous-discrete data representation comes with the challenge of joint optimization. An alternative is to model a unified continuous data representation that circumvents these challenges. StaSy [24] utilizes a naive one-hot encoding representation to unify the data space but necessitates additional computationally heavy tricks to improve its performance. It trains a diffusion model on a subset of data before gradually extending to the whole training set. Thus, in addition to the diffusion model parameter, $\theta$, the model is required to learn a selection importance vector, $\mathbf{v}\in[0,1]$ through an alternative convex search [2]. The search solves the loss function by iteratively optimizing between $\theta$ and $\mathbf{v}$ while also requiring a self-paced regularizer, $r(.)$, that requires fine-tuning. Another tabular diffusion method, TabSYN [58], unifies the data space by transforming discrete features using a learned $\beta$-VAE latent embedding. The architecture of the VAE consists of a tokenizer that encodes discrete features into a one-hot vector, then parameterizes each category as a learnable vector learned by a Transformer [52] encoder. Likewise to StaSy, TabSYN is also computationally expensive since it relies on training a separate VAE before the diffusion process to obtain latent codes. Additionally, the quality of synthetic data is highly dependent on the quality of the latent space. In the following section, we show that leveraging our simple and effective continuous data representation to train tabular diffusion models is essential in synthesizing high-quality tabular data.
4 Method
Tabular data undergo preprocessing preceding the diffusion generative process, and transforming them into intelligible representations streamlines and improves the training of diffusion models.
4.1 Geometric Implications on the Data Representation
In traditional deep learning [16], a sparse representation suffers from the curse of dimensionality [3], where the feature space grows exponentially with the number of categories, reducing model generalization [27, 41]. While maintaining a dense representation is key, representation learning also [4, 29] emphasizes the importance of separability, enabling the neural network to learn decision boundaries in the continuous embedding space. In the following excerpt, we find that balancing both density and separability while incorporating order is essential for unified tabular diffusion.
Density. Diffusion models, which learn to generate data through iterative perturbations and reconstructions of noise, encounter geometric challenges when applied to high-dimensional, sparse representations of categorical features. In this discussion, we use the sparse one-hot representation as an exemplar to provide geometric insights into these challenges.
Let $\{e_{1},e_{2},\dots,e_{K}\}\subset\mathbb{R}^{K}$ denote the set of one-hot vectors, where
$e_{k}=(0,\dots,0,\underbrace{1}_{k\text{-th entry}},0,\dots,0),$ (6)
represents the $k$-th category. For any point $x\in\mathbb{R}^{K}$, let
$d_{k}(x)=\|x-e_{k}\|_{2},$ (7)
denote the Euclidean distance between $x$ and $e_{k}$. On the data manifold, there exist “singular” [55] points where the vector field is hard to learn for diffusion models with Gaussian transitions. We demonstrate this with a uniform $K$-categorical distribution. Specifically, we define:
Definition 4.1 ($n$-singular point).
A point $x\in\mathbb{R}^{K}$ is an $n$-singular point if there exists a subset $\mathcal{S}\subseteq\{1,\dots,K\}$ with $|\mathcal{S}|=n$ such that:
  1. 1.
    $d_{k}(x)=d_{k^{\prime}}(x)$,  $\forall k,k^{\prime}\in\mathcal{S}$.
  2. 2.
    $d_{k}(x)\neq d_{m}(x)$,  $\forall k\in\mathcal{S}$, $\forall m\notin\mathcal{S}$.
An $n$-singular point can be extended as a minimal $n$-singular point if it satisfies Definition 4.1, and minimizes the Euclidean distance to all one-hot vectors in $\mathcal{S}$: $\min_{x}\|x-e_{k}\|_{2}$ $\forall k\in\mathcal{S}$. Hence, the minimal $n$-singular point is given by:
$x_{\mathcal{S}}^{(n)}=\frac{1}{n}\sum_{k\in\mathcal{S}}e_{k},$ (8)
that corresponds to the centroid of the $n$ one-hot vectors.
Definition 4.2 ($n$-singular hyperplane).
For each minimal $n$-singular point, there exists an $n$-singular hyperplane that is comprised of the set of all $n$-singular points associated with its respective $n$ one-hot vectors in $\mathcal{S}$. Formally, it is defined as:
$\!\!\!H_{\mathcal{S}}\;=\;\bigl{\{}x\in\mathbb{R}^{K}\;\big{|}\;d_{k}(x)=d_{k^{\prime}}(x),\quad\forall k,k^{\prime}\in\mathcal{S}\bigr{\}},$ (9)
where $H_{\mathcal{S}}$ is an affine subspace in $\mathbb{R}^{K}$ of dimension $\dim(H_{\mathcal{S}})=K-|\mathcal{S}|+1$. The hyperplane spans the corresponding minimal $n$-singular point $x_{\mathcal{S}}^{(n)}$ and non-minimal $n$-singular points. For each $n<K$, there are $\binom{K}{n}$ distinct $n$-singular hyperplanes, one for each subset $\mathcal{S}$ of size $n$. Across all $2\leq n\leq K$, the number of minimal $n$-singular points on the probability simplex scales combinatorially: $\sum_{n=2}^{K}\binom{K}{n}=2^{K}-(K+1)$. Thus, each minimal singular point carries the additional complexity of a continuous singular hyperplane.
Diffusion models rely on gradients derived from a learned vector field to denoise data iteratively [20, 31]. For regions within proximity of the singular hyperplanes, learning the gradients of diffusion models suffers from high variance due to conflicting directions arising from equidistant one-hot points. We show that the variance of the conditional score function increases asymptotically with the degree of $n$-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$. We derive the variance of the conditional score function evaluated at a minimal $n$-singular point as:
$\mathrm{Var}(g|x)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\frac{n-1}{n},$ (10)
Figure 2: Singular Regions in a 3D One-Hot Setting.
where we define the conditional score and expected score as $g$ and $\bar{g}$. See Appendix A.1 for a proof. We find that near a minimal $n$-singular point, $x$, the posterior-weighted variance of the conditional score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$, leading the score variance to approach zero.
In Figure 2, we depict a three-dimensional setting of the one-hot representation. As illustrated, a minimal $3$-singular point occurs at the (red) centroid, $\bigl{(}\frac{1}{3},\frac{1}{3},\frac{1}{3}\bigr{)}$. Along with the non-minimal $3$-singular points, these points form a (red-dashed) $1$-dimensional singular hyperplane $H_{\{1,2,3\}}$ perpendicular to the centroid of the probability simplex formed by $e_{1},e_{2},e_{3}$. Similarly, there are $\binom{3}{2}$ $2$-singular (blue) points accompanied by their respective (blue-dashed) hyperplanes $H_{\{1,2\}}$, $H_{\{1,3\}}$, and $H_{\{2,3\}}$, each of which is a $2$-dimensional affine subspace (plane). These singular hyperplanes pose more difficulty for diffusion models to learn effectively, especially when $K$ is large.
Separability and Order. A sparse representation naturally accommodates separability for nominal features, enabling one-hot to assign each category to each dimension. However, higher-dimensional spaces introduce higher-order singular hyperplanes that complicate training. This presents a density-separability trade-off. In our experiments (Table 3), we discover that sparse representations has a significant impact on harming diffusion generative performance. Thus, our initial goal is to reduce the dimensionality when designing our representation.
Methods like Analog Bits [10] admit a similar idea of shrinking data dimension. Hence, a potential concern of a dense representation is to retain their ability of representing nominal features. However, the notion of separability [5] demonstrates that nominal features can still be effectively encoded by dense representations, provided they are sufficiently separated within the embedding space. This perspective aligns with learned entity embeddings [17], which demonstrate that nominal categories—despite lacking inherent order—can be effectively represented as low-dimensional embeddings. However, for datasets with a large presence of ordinal features such as “Education” in Adult and “Day” in Beijing, our experiment in Appendix B.2, Table 7 highlights that order is an important factor for ordinal features. Therefore, designing a dense representation that is separable and capable of encoding ordinal structure is critical for encoding both nominal and ordinal features.
4.2 TabRep Architecture
CatConverter. Inspired by the discrete Fourier transform (DFT) [39, 6], we draw on the concept of roots of unity to design a continuous representation for diffusion models to generate categorical variables. We refer to our representation as the CatConverter. In harmonic analysis, the DFT maps signals into the frequency domain using complex exponentials, where the $K$-th roots of unity represent equally spaced points on the unit circle, given by phases:
$\theta_{k}=\frac{2\pi k}{K_{j}^{(i)}}\quad\text{for }k=0,1,\dots,K_{j}^{(i)}-1.$ (11)
Analogously, we treat a categorical feature $c_{j}^{(i)}$ with $K_{j}^{(i)}$ distinct values as selecting one of these $K_{j}^{(i)}$ points on the unit circle. Each category is thus mapped to a unique phase, and we represent it using the real and imaginary components of the corresponding complex exponential:
$\text{CatConverter}(c_{j}^{(i)},K_{j}^{(i)})=\left[\cos\left(\frac{2\pi c_{j}^{(i)}}{K_{j}^{(i)}}\right),\sin\left(\frac{2\pi c_{j}^{(i)}}{K_{j}^{(i)}}\right)\right],$ (12)
Table 1: Categorical Representation Dimensions.
Representation Dimensions
One-Hot $\sum_{j=1}^{D_{\text{cat}}}K_{j}$
Learned Embedding $d_{\text{emb}}\cdot D_{\text{cat}}$
Analog Bits $\sum_{j=1}^{D_{\text{cat}}}\lceil\log_{2}(K_{j})\rceil$
Dictionary $D_{\text{cat}}$
CatConverter $2\cdot D_{\text{cat}}$
where $\text{CatConverter}()\in\mathbb{R}^{2\cdot D_{\text{CAT}}}$. Viewing a categorical entry as a discrete harmonic index enables us to embed the category in a two-dimensional phase space. This viewpoint retains the geometric insight from the DFT—the roots of unity form a symmetric and uniform structure on the unit circle—providing a smooth, dense, and geometry-aware representation for categorical variables.
In Table 1, CatConverter offers a dense 2D representation relative to alternate categorical representations. For CatConverter, there exists one minimal $K$-singular point. Despite that, and $K$ number of $2$-singular points, there are no $n$ singular points or hyperplanes for $2<n<K$. Therefore, any $n$-singular point for $n>2$ coincides with the minimal $K$-singular point in this 2D representation.
Figure 3: Separability of CatConverter. CatConverter preserves nominal features for up to $128$ categories.
Next, we show that CatConverter’s representation has ample separability for handling high-cardinality nominal categorical features commonly found in tabular datasets. As illustrated in Figure 3, our representation could be easily distinguished by a small MLP among $128$ categories mapped onto a continuous phase space.
In addition to nominal features, the separability of CatConverter coupled with its circular geometry naturally accommodates both periodic and ordinal features. This facilitates an enhanced preservation of the feature’s intrinsic nature and characteristics. Note that prior to CatConverter, we introduced a static one-dimensional embedding, Dictionary (dic), but we found that it underperformed.
Algorithm 1 Training TabRep-DDPM/Flow
1:DDPM:                                                          Flow Matching:
2:while not converged do
3:   Sample $\mathbf{z}_{0}=[\mathbf{x}_{0},\mathbf{c}_{0}]\sim p(\mathbf{z})$
4:   Encode $\mathbf{c}_{0}^{\textsc{CC}}\leftarrow\text{CatConverter}(\mathbf{c}_{0},\mathbf{K})$
5:   Encode $\mathbf{x}_{0}\leftarrow\textsc{QuantileTransformer}(\mathbf{x}_{0})$
6:   $\mathbf{z}_{0}^{\textsc{CC}}\leftarrow\text{CONCAT}(\mathbf{x}_{0},\mathbf{c}_{0}^{\textsc{CC}})$
7:   Sample $t\sim\text{Uniform}(\{1,\ldots,T\})$
8:   Sample noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
9:    {tightgraybox_train}
10:   Compute $\mathbf{z}_{t}=\alpha_{t}^{\mathbf{z}}\mathbf{z}_{0}^{\textsc{CC}}+\sigma_{t}^{\mathbf{z}}\bm{\epsilon}$                               Compute $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}^{\textsc{CC}}+t\bm{\epsilon}$
11:                                                                    Define $\mathbf{u}_{t}(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})=\bm{\epsilon}-\mathbf{z}_{0}^{\textsc{CC}}$
12:   Compute $\mathcal{L}_{\textsc{TabRep-DDPM}}=$                                  Compute $\mathcal{L}_{\textsc{TabRep-Flow}}=$
13:   $\mathbb{E}_{t,\mathbf{z}_{0}^{\textsc{CC}},\mathbf{z}_{t}}\left[\left\lVert\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})-\nabla_{\mathbf{z}_{t}}\log p_{\theta}(\mathbf{z}_{t})\right\rVert^{2}\right]$     $\mathbb{E}_{t,\mathbf{z}_{0}^{\textsc{CC}},\mathbf{z}_{t}}\left[\left\lVert\mathbf{v}_{\theta}(\mathbf{z}_{t})-\mathbf{u}_{t}(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})\right\rVert^{2}\right]$
14:   Update $\theta$ via gradient descent for $\nabla_{\theta}\mathcal{L}_{\textsc{TabRep-DDPM}}$ and $\nabla_{\theta}\mathcal{L}_{\textsc{TabRep-Flow}}$
15:end while
Algorithm 2 Sampling TabRep-DDPM/Flow
1:DDPM:                                                          Flow Matching:
2:Sample $\mathbf{z}_{T}^{\text{CC}}\sim\mathcal{N}(0,\mathbf{I})$
3:for $t=T,\ldots,1$ do
4:    {tightgraybox_sample}
5:   Sample $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$,                              Discretize time $t_{i}=i/T$,
6:   else $\bm{\epsilon}=0$                                                     for $i=T,T-1,\ldots,1$
7:   $\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}^{\mathbf{z}}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}^{\mathbf{z}}}{\sqrt{1-\bar{\alpha}_{t}^{\mathbf{z}}}}\nabla_{\mathbf{z}_{t}}\log p_{\theta}(\mathbf{z}_{t})\right)+\sigma_{t}^{\mathbf{z}}\bm{\epsilon}$    $\mathbf{z}_{t_{i}-1}=\mathbf{z}_{t_{i}}-\frac{1}{T}\cdot\mathbf{v_{\theta}}(\mathbf{z}_{t_{i}})$,
8:                                                                     via Euler ODE Solver
9:end for
10:Output $\mathbf{z}_{0}^{\textsc{CC}}=[\mathbf{x}_{0},\mathbf{c}_{0}^{\textsc{CC}}]$
11:Decode $\mathbf{c}_{0}\leftarrow\text{InvCatConverter}(\mathbf{c}_{0}^{\textsc{CC}},\mathbf{K})$
12:Decode $\mathbf{x}_{0}\leftarrow\textsc{InvQuantileTransformer}(\mathbf{x}_{0})$
13:$\mathbf{z}_{0}\leftarrow\text{CONCAT}(\mathbf{x}_{0},\mathbf{c}_{0})$
14:return $\mathbf{z}_{0}^{\textsc{CC}}$
Figure 4: Training and sampling algorithms of TabRep-DDPM/Flow.
Diffusion Model. We demonstrate that TabRep’s representation is effective when modeled by either a DDPM [19] or a Flow Matching [31] unified continuous diffusion process. To unify the dataspace, we represent discrete variables via $\text{CatConverter}(\mathbf{c}_{0},\mathbf{K})$ and concatenate with continuous features forming our dataset. Our dataset can be denoted as $\{\mathbf{z}_{0}^{\textsc{CC}}\}=\{[\mathbf{x}_{0},\mathbf{c}_{0}^{\textsc{CC}}]\}$.
For DDPM [19], we define the forward process by progressively perturbing the data distribution using a Gaussian noise model, where the latent state at time $t$, $\mathbf{z}_{t}$, is computed as $\mathbf{z}_{t}=\alpha_{t}^{\mathbf{z}}\mathbf{z}_{0}^{\textsc{CC}}+\sigma_{t}^{\mathbf{z}}\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The denoising model $p_{\theta}(\mathbf{z}_{t})$ is trained to predict the posterior gradients $\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})$, minimizing the weighted variance loss $\mathcal{L}_{\textsc{TabRep-DDPM}}$:
$\mathbb{E}_{t,\mathbf{z}_{0}^{\textsc{CC}},\mathbf{z}_{t}}\left[\left\lVert\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})-\nabla_{\mathbf{z}_{t}}\log p_{\theta}(\mathbf{z}_{t})\right\rVert^{2}\right].$ (13)
For Flow Matching [31], we instead define the dynamics in terms of a conditional vector field $\mathbf{u}_{t}(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})=\bm{\epsilon}-\mathbf{z}_{0}^{\textsc{CC}}$ with $\mathbf{z}_{t}:=(1-t)\mathbf{z}_{0}^{\textsc{CC}}+t\bm{\epsilon}$. The model learns the target field by minimizing the discrepancy between the predicted vector $\mathbf{v}_{\theta}(\mathbf{z}_{t})$ and the ground truth $\mathbf{u}_{t}(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})$ through the flow matching loss $\mathcal{L}_{\textsc{TabRep-Flow}}$:
$\mathbb{E}_{t,\mathbf{z}_{0}^{\textsc{CC}},\mathbf{z}_{t}}\left[\left\lVert\mathbf{v}_{\theta}(\mathbf{z}_{t})-\mathbf{u}_{t}(\mathbf{z}_{t}|\mathbf{z}_{0}^{\textsc{CC}})\right\rVert^{2}\right].$ (14)
At sampling time, TabRep-DDPM performs reverse diffusion by iteratively denoising $\mathbf{z}_{t}$ back to $\mathbf{z}_{0}^{\textsc{CC}}$, while TabRep-Flow solves a deterministic ordinary differential equation (ODE) of the vector field. Our complete training and sampling algorithms can be found in Figure 4.
5 Experiments
We evaluate the performance of TabRep-DDPM and TabRep-Flow. Our results are compared against various datasets, baselines, and benchmarks.
5.1 Setup
Implementation Information. Experimental results are obtained over an average of $20$ sampling seeds using the best-validated model. Continuous features are encoded using a QuantileTransformer [40]. The order in which the categories of a feature are assigned in our encoding schemes is based on lexicographic ordering for simplicity. Further details are in Appendix B.
Datasets. We select seven datasets from the UCI Machine Learning Repository to conduct our experiments. This includes Adult, Default, Shoppers, Stroke, Diabetes, Beijing, and News which contain a mix of continuous and discrete features. Further details are in Appendix C.1.
Baselines. We compare our model against existing diffusion baselines for tabular generation since they are the best-performing. This includes STaSy [24], CoDi [30], TabDDPM [26], TabSYN [58], and TabDiff [47]. Further details are in Appendix C.2.
Benchmarks. We observe that downstream task performance measured by machine learning efficiency (MLE) typically translates to most other benchmarks. Thus, the primary quality benchmark in our main paper will be MLE for brevity, and a privacy benchmark, membership inference attacks (MIA). The remaining fidelity benchmarks include: column-wise density (CWD), pairwise-column correlation (PCC), $\alpha$-precision, $\beta$-recall, and classifier-two-sample test (C2ST). Further details regarding benchmark information and additional results are in Appendix C.3 and D.
5.2 Ablation Studies
Table 2: Ablation study on TabRep’s unified representation versus a separate representation.
AUC $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$
Methods Adult Default Shoppers Stroke Diabetes Beijing News
TabDDPM 0.910$\pm$.001 0.761$\pm$.004 0.915$\pm$.004 0.808$\pm$.033 0.376$\pm$.003 0.592$\pm$.012 3.46$\pm$1.25
TabRep-DDPM 0.913$\pm$.002 0.764$\pm$.005 0.926$\pm$.005 0.869$\pm$.027 0.373$\pm$.003 0.508$\pm$.006 0.836$\pm$.001
TabFlow 0.908$\pm$.002 0.742$\pm$.008 0.914$\pm$.005 0.821$\pm$.082 0.377$\pm$.002 0.574$\pm$.010 0.850$\pm$.017
TabRep-Flow 0.912$\pm$.002 0.782$\pm$.005 0.919$\pm$.005 0.830$\pm$.028 0.377$\pm$.002 0.536$\pm$.006 0.814$\pm$.002
Table 3: Ablation study on categorical representations under a unified continuous data space.
AUC $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$
Methods Adult Default Shoppers Stroke Diabetes Beijing News
OneHot-DDPM 0.476$\pm$.057 0.557$\pm$.052 0.799$\pm$.126 0.797$\pm$.133 0.363$\pm$.008 2.143$\pm$.339 0.840$\pm$.020
Learned1D-DDPM 0.611$\pm$.008 0.575$\pm$.012 0.876$\pm$.028 0.743$\pm$.032 0.179$\pm$.010 0.921$\pm$.006 0.858$\pm$.010
Learned2D-DDPM 0.793$\pm$.003 0.290$\pm$.009 0.103$\pm$.011 0.850$\pm$.035 0.205$\pm$.007 0.969$\pm$.005 0.857$\pm$.023
i2b-DDPM 0.911$\pm$.001 0.762$\pm$.003 0.919$\pm$.004 0.852$\pm$.029 0.370$\pm$.008 0.542$\pm$.008 0.844$\pm$.013
dic-DDPM 0.912$\pm$.002 0.763$\pm$.003 0.910$\pm$.004 0.824$\pm$.027 0.375$\pm$.006 0.547$\pm$.008 0.851$\pm$.013
TabRep-DDPM 0.913$\pm$.002 0.764$\pm$.005 0.926$\pm$.005 0.869$\pm$.027 0.373$\pm$.003 0.508$\pm$.006 0.836$\pm$.001
OneHot-Flow 0.895$\pm$.003 0.759$\pm$.005 0.910$\pm$.006 0.812$\pm$.129 0.372$\pm$.005 0.765$\pm$.016 0.850$\pm$.017
Learned1D-Flow 0.260$\pm$.014 0.438$\pm$.009 0.134$\pm$.015 0.142$\pm$.033 0.184$\pm$.007 0.806$\pm$.009 0.873$\pm$.007
Learned2D-Flow 0.126$\pm$.012 0.709$\pm$.008 0.868$\pm$.007 0.180$\pm$.030 0.177$\pm$.008 0.787$\pm$.007 0.866$\pm$.005
i2b-Flow 0.911$\pm$.001 0.763$\pm$.004 0.910$\pm$.005 0.797$\pm$.027 0.372$\pm$.003 0.543$\pm$.007 0.847$\pm$.014
dic-Flow 0.912$\pm$.002 0.763$\pm$.004 0.903$\pm$.005 0.807$\pm$.028 0.376$\pm$.007 0.561$\pm$.007 0.853$\pm$.014
TabRep-Flow 0.912$\pm$.002 0.782$\pm$.005 0.919$\pm$.005 0.830$\pm$.028 0.377$\pm$.002 0.536$\pm$.006 0.814$\pm$.002
Unified and Separate Data Representations. We analyze the impact of training tabular diffusion models between our unified data representation and a separate data representation. In Table 2, we compare TabRep-DDPM and TabRep-Flow to TabDDPM [26] and TabFlow. Note that we introduce TabFlow, by modeling continuous features using Flow Matching [31, 33] and categorical features using Discrete Flow Matching [8] under the same separate continuous-discrete data representation as TabDDPM. Results illustrate that data synthesized by diffusion-based models using TabRep’s unified representation consistently outperforms a separate representation when training diffusion and flow matching models.
Categorical Representations. We conduct ablation studies with respect to categorical data representations used in existing diffusion baselines to assess the performance of diffusion models under a unified data representation. This includes one-hot encoding used in [24, 30, 47], learned one-dimensional and two-dimensional embeddings [34, 17], and Analog Bits (i2b) [10] from the discrete image diffusion domain. Additionally, we introduce an intuitive static one-dimensional embedding, Dictionary (dic). Results in Table 3 showcase that TabRep outperforms the ablated representations in generating data under a unified data representation via diffusion models. Further details regarding the various representations are in Appendix B.1.
5.3 Baseline Performance
Generation Quality. We benchmark TabRep against baselines on a downstream MLE task where we determine the AUC and RMSE of XGBoost [9] for classification and regression tasks on the generated synthetic datasets. The diffusion models are trained on the training set, and synthetic samples of equivalent size are generated, which are then evaluated using the testing set. As observed in Table 4, CatConverter is effective on both DDPM and Flow Matching since TabRep-DDPM and TabRep-Flow consistently attain the best MLE performance compared to existing baselines. Additionally, our method is the first to yield performance levels greater than the Real datasets including Default, Stroke, and News.
Table 4: AUC (classification) and RMSE (regression) scores of Machine Learning Efficiency. Higher scores indicate better performance.
AUC $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$
Methods Adult Default Shoppers Stroke Diabetes Beijing News
Real 0.927$\pm$.000 0.770$\pm$.005 0.926$\pm$.001 0.852$\pm$.002 0.384$\pm$.003 0.423$\pm$.003 0.842$\pm$.002
STaSy 0.906$\pm$.001 0.752$\pm$.006 0.914$\pm$.005 0.833$\pm$.030 0.374$\pm$.003 0.656$\pm$.014 0.871$\pm$.002
CoDi 0.871$\pm$.006 0.525$\pm$.006 0.865$\pm$.006 0.798$\pm$.032 0.288$\pm$.009 0.818$\pm$.021 1.21$\pm$.005
TabDDPM 0.910$\pm$.001 0.761$\pm$.004 0.915$\pm$.004 0.808$\pm$.033 0.376$\pm$.003 0.592$\pm$.012 3.46$\pm$1.25
TabSYN 0.906$\pm$.001 0.755$\pm$.004 0.918$\pm$.004 0.845$\pm$.035 0.361$\pm$.001 0.586$\pm$.013 0.862$\pm$.021
TabDiff 0.912$\pm$.002 0.763$\pm$.005 0.919$\pm$.005 0.848$\pm$.021 0.353$\pm$.006 0.565$\pm$.011 0.866$\pm$.021
TabRep-DDPM 0.913$\pm$.002 0.764$\pm$.005 0.926$\pm$.005 0.869$\pm$.027 0.373$\pm$.003 0.508$\pm$.006 0.836$\pm$.001
TabRep-Flow 0.912$\pm$.002 0.782$\pm$.005 0.919$\pm$.005 0.854$\pm$.028 0.377$\pm$.002 0.536$\pm$.006 0.814$\pm$.002
Table 5: Recall Scores of MIAs. A score closer to $50\%$ is better for privacy-preservation.
Methods Adult Default Shoppers Stroke Diabetes Beijing News
STaSy 24.51$\pm$0.44 30.37$\pm$0.99 17.54$\pm$0.19 34.63$\pm$1.39 29.75$\pm$0.16 34.06$\pm$0.33 23.49$\pm$0.67
CoDi 0.05$\pm$0.01 3.41$\pm$0.53 1.36$\pm$0.45 27.32$\pm$3.50 0.00$\pm$0.00 0.40$\pm$0.09 0.04$\pm$0.02
TabDDPM 56.90$\pm$0.29 44.96$\pm$0.59 46.08$\pm$1.57 55.12$\pm$0.81 52.06$\pm$0.11 48.35$\pm$0.79 9.88$\pm$0.52
TabSYN 42.91$\pm$0.31 43.71$\pm$0.89 42.14$\pm$0.76 47.97$\pm$1.27 44.42$\pm$0.28 46.53$\pm$0.52 34.42$\pm$0.90
TabDiff 52.00$\pm$0.35 46.67$\pm$0.55 46.86$\pm$1.20 47.15$\pm$1.30 31.01$\pm$0.22 48.20$\pm$0.65 16.03$\pm$0.66
TabRep-DDPM 52.78$\pm$0.21 48.96$\pm$0.41 48.16$\pm$0.90 50.89$\pm$0.93 51.43$\pm$0.13 49.50$\pm$0.52 40.10$\pm$0.55
TabRep-Flow 50.51$\pm$0.21 47.07$\pm$0.40 49.19$\pm$0.86 49.43$\pm$1.16 50.33$\pm$0.13 49.14$\pm$0.81 35.48$\pm$0.78
Privacy Preservation. We perform membership inference attacks (MIAs) to evaluate privacy preservation by assessing the vulnerability of the methods to privacy leakage [48]. A privacy-preserving model yields recall scores of 50%, indicating that the attack is no better than random guessing. In Table 5, our method preserves privacy with MIAs scoring close to 50% for recall. Precision scores can be found in Appendix D.1 which delivers similar privacy preservation insights.
Table 6: Training and Sampling Duration.
Methods Training (s) Sampling (s) Total (s)
STaSy 6608.84 12.93 6621.77
CoDi 24039.96 9.41 24049.37
TabDDPM 3112.07 66.82 3178.89
TabSYN 2373.98 $+$ 1084.82 10.54 3469.30
TabDiff 5640 15.2 5655.2
TabRep-DDPM 2070.59 59.00 2130.59
TabRep-Flow 2028.33 3.07 2031.40
Generation Efficacy. Training and Sampling TabRep is the most efficient among all baselines and does not necessitate additional computing power. We compare the generation efficacy using the Adult dataset with a balanced mix of continuous and discrete features. In Table 6, we observe that our method is the quickest to train and sample in duration.
(a) AUC vs. Train Epochs
(b) AUC vs. Sample NFEs
Figure 5: Training and Sampling Efficacy.
We also conducted experiments on the convergence speed for the training process. As illustrated in Figure 5, our method converges to a high AUC earliest in the training stages. Lastly, we assess the number of function evaluations (NFEs) the models take to sample. Figure 5 indicates that TabRep-Flow can attain its best performance as early as $8$ NFEs. At $1000$ NFEs, we observe that both TabRep models are the best-performing.
6 Conclusion
In this work, we present TabRep, a simple and effective continuous representation for training tabular diffusion models. Motivated by geometric implications on the data manifold, our representation is dense, separable, and captures meaningful information for diffusion models. We conducted extensive experiments on a myriad of datasets, baselines, and benchmarks to evaluate TabRep. The results showcase TabRep’s prowess in generating high-quality, privacy-preserving synthetic data while remaining computationally inexpensive.
Appendix
[…]
Appendix A Proofs
A.1 Variance of Learning Gradients in Diffusion Models
We show that the variance of the score function at a minimal $n$-singular point increases asymptotically with respect to $n$ dimensions. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as:
$p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$ (15)
Hence, the score function of the conditional distribution $p_{t}(x|e_{k})$ is given by:
$\nabla_{x}\log p_{t}(x|e_{k})$ $=\nabla_{x}\log\left[\frac{1}{(2\pi\sigma_{t}^{2})^{K/2}}\exp\left(-\frac{1}{2\sigma_{t}^{2}}||x-\alpha_{t}e_{k}||^{2}\right)\right]$ (16)
$=\nabla_{x}\left[-\frac{K}{2}\log(2\pi\sigma_{t}^{2})-\frac{1}{2\sigma^{2}_{t}}||x-\alpha_{t}e_{k}||^{2}\right]$ (17)
$=-\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}e_{k})$ (18)
In which we define it as the gradient:
$g_{k}(x):=\nabla_{x}\log p_{t}(x|e_{k})=-\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}e_{k})$ (19)
Assume we have a uniform prior over categories:
$p(e_{k})=\frac{1}{K}\quad\forall k\in\mathcal{S}=\{1,\ldots,K\}$ (20)
Per Bayes’ Rule, we compute the posterior over $e_{k}$:
$q(e_{k}|x)$ $=\frac{p(x|e_{k})p(e_{k})}{\sum_{m=1}^{K}p(x|e_{m})p(e_{m})}$ (21)
$=\frac{p(x|e_{k})}{\sum_{m=1}^{K}p(x|e_{m})}$ (22)
$=\frac{p(x|e_{k})}{\sum_{k\in S}p(x|e_{k})+\sum_{m\notin S}p(x|e_{m})}$ (23)
Since the forward process is modeled as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$, we can infer that:
$p(x|e_{k})\propto\exp\left(-\frac{1}{2\sigma_{t}^{2}}||x-\alpha_{t}e_{k}||^{2}\right)$ (24)
Then, $\forall k\in\mathcal{S}$, the likelihood terms are identical:
$p_{t}(x|e_{k})=\exp\left(-\frac{d_{k}^{2}}{2\sigma_{t}^{2}}\right):=A$ (25)
where $d$ is defined in Equation 7. $\forall m\notin S$, $d_{m}(x)^{2}>d_{k}(x)^{2}$, thus:
$p_{t}(x|e_{m})=\exp\left(-\frac{d_{m}^{2}}{2\sigma_{t}^{2}}\right)\ll A$ (26)
Therefore, the posterior simplifies to:
$q(e_{k}|x)$ $=\frac{p(x|e_{k})}{\sum_{k\in S}p(x|e_{k})}$ (27)
$\approx\frac{A}{nA}$ (28)
$=\frac{1}{n},\quad\forall k\in\mathcal{S}.$ (29)
We now compute the expected score:
$\bar{g}(x)$ $=\sum_{k=1}^{K}q(e_{k}|x)\cdot g_{k}(x)$ (30)
$\approx\sum_{k\in\mathcal{S}}\frac{1}{n}\cdot g_{k}(x)$ (31)
$=-\frac{1}{n\sigma_{t}^{2}}\sum_{k\in\mathcal{S}}(x-\alpha_{t}e_{k})$ (32)
$=-\frac{1}{\sigma_{t}^{2}}\left(x-\alpha_{t}\bar{e}\right),$ (33)
where $\bar{e}:=\frac{1}{n}\sum_{k\in\mathcal{S}}e_{k}$ is the centroid of the vectors in $\mathcal{S}$. Next, compute the variance:
$\mathrm{Var}(g|x)$ $=\sum_{k\in\mathcal{S}}\frac{1}{n}\left\|g_{k}(x)-\bar{g}(x)\right\|^{2}$ (34)
$=\sum_{k\in\mathcal{S}}\frac{1}{n}\left\|-\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}e_{k})+\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}\bar{e})\right\|^{2}$ (35)
$=\sum_{k\in\mathcal{S}}\frac{1}{n}\cdot\frac{1}{\sigma_{t}^{4}}\left\|\alpha_{t}(\bar{e}-e_{k})\right\|^{2}$ (36)
$=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\sum_{k\in\mathcal{S}}\frac{1}{n}\left\|\bar{e}-e_{k}\right\|^{2}$ (37)
To complete the variance computation, we compute the squared distance between each one-hot vector and the centroid:
$\left\|\bar{e}-e_{k}\right\|^{2}$ $=\sum_{i=1}^{K}\left(\bar{e}_{i}-e_{k}(i)\right)^{2}$ (38)
$=\left(\frac{n-1}{n}\right)^{2}+(n-1)\cdot\left(\frac{1}{n}\right)^{2}$ (39)
$=\frac{n-1}{n}$ (40)
Note that for $i=k\in S$, $\bar{e}_{k}=\frac{1}{n},e_{k}(k)=1$ and for $i\neq k\in S$, $e_{k}(k)=0$. Substituting into the variance:
$\mathrm{Var}(g|x)$ $=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\cdot\sum_{k\in\mathcal{S}}\frac{1}{n}\cdot\frac{n-1}{n}$ (41)
$=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\cdot\frac{n-1}{n}$ (42)
We find that at a minimal $n$-singular point $x$, the posterior-weighted variance of the score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$:
$q(e_{k^{*}}|x)\approx 1,\quad q(e_{k}|x)\approx 0\quad\forall k\neq k^{*}$ (43)
Then, the expected score becomes:
$\bar{g}(x)=\sum_{k=1}^{K}q(e_{k}|x)\cdot g_{k}(x)\approx g_{k^{*}}(x)$ (44)
And the variance reduces to:
$\mathrm{Var}(g|x)$ $=\sum_{k=1}^{K}q(e_{k}|x)\cdot\left\|g_{k}(x)-\bar{g}(x)\right\|^{2}$ (45)
$\approx 1\cdot\left\|g_{k^{*}}(x)-g_{k^{*}}(x)\right\|^{2}+\sum_{k\neq k^{*}}0\cdot\left\|g_{k}(x)-g_{k^{*}}(x)\right\|^{2}$ (46)
$=0$ (47)
We now generalize the above result to an arbitrary categorical prior $\Pi=\{\pi_{k}\}_{k\in\mathcal{S}}$ over categories $\{e_{k}\}_{k\in\mathcal{S}}$, where $x_{0}\sim\Pi$ and the forward diffusion process is defined as:
$p_{t}(x|x_{0})=\mathcal{N}(x|\alpha_{t}x_{0},\sigma_{t}^{2}I).$ (48)
The marginal likelihood is then given by:
$p_{t}(x)=\sum_{k\in\mathcal{S}}p_{t}(x|x_{0}=e_{k})\pi_{k}=\sum_{k\in\mathcal{S}}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\pi_{k}.$ (49)
Suppose we observe a noised sample $x\in\mathbb{R}^{k}$ at time $t$. Then the posterior probability that $x$ was generated by adding noise to $e_{k}$ is:
$q_{t}(x_{0}=e_{k}|x)=\frac{p_{t}(x|x_{0}=e_{k})\pi_{k}}{\sum_{m\in\mathcal{S}}p_{t}(x|x_{0}=e_{j})\pi_{m}}.$ (50)
We compute the expected score at time $t$ as follows:
$\mathbb{E}_{p_{t}(x)}[\nabla_{x}\log p_{t}(x)]$ $=\mathbb{E}_{x_{0}\sim\Pi}\left[\mathbb{E}_{p_{t}(x|x_{0})}\left[\nabla_{x}\log p_{t}(x|x_{0})\right]\right]$ (51)
$=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\nabla_{x}\log\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I).$ (52)
The score of the Gaussian is:
$\nabla_{x}\log\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)=-\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}e_{k})=\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}.$ (53)
Therefore, the expected score becomes:
$\bar{g}:=\mathbb{E}_{p_{t}(x)}[\nabla_{x}\log p_{t}(x)]$ $=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\cdot\left(\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}\right)$ (54)
$=C\sum_{k\in\mathcal{S}}\pi_{k}\exp\left(-\frac{\|x-\alpha_{t}e_{k}\|^{2}}{2\sigma_{t}^{2}}\right)\left(\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}\right),$ (55)
where $C:=\frac{1}{(2\pi)^{k/2}\sigma_{t}^{k}}$.
Next, we compute the variance of the conditional score around its expectation:
$\mathbb{E}_{x_{0}\sim\Pi}\left[\mathbb{E}_{p_{t}(x|x_{0})}\left[\left\|\nabla_{x}\log p_{t}(x|x_{0})-\bar{g}\right\|^{2}\right]\right]$ $=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\left\|\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}-\bar{g}\right\|^{2}$ (56)
$=C\sum_{k\in\mathcal{S}}\pi_{k}\exp\left(-\frac{\|x-\alpha_{t}e_{k}\|^{2}}{2\sigma_{t}^{2}}\right)\left\|\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}-\bar{g}\right\|^{2}.$ (57)
This formulation reveals that score variance is low when the posterior is sharply peaked (i.e., $x$ is close to a single one-hot vector), and increases when the posterior mass is spread over multiple categories. The minimal $n$-singular point case is recovered when $\pi_{k}=\frac{1}{n}$ for $k\in\mathcal{S}$ and $x=\frac{1}{n}\sum_{k\in\mathcal{S}}e_{k}$, confirming the consistency of both analyses.
Appendix B Implementation
The following delineates the foundation of our experiments:
  • Codebase: Python & PyTorch
  • CPU: AMD EPYC-Rome 7002
  • GPU: NVIDIA A100 80GB PCIe
B.1 Categorical Representations
One-Hot Encoding. The one-hot encoding (OneHot) representation of $\mathbf{c}^{(i)}$ is constructed by concatenating the one-hot encoded vectors of each individual feature $c_{j}^{(i)}$. Specifically, the one-hot encoding for $c_{j}^{(i)}$ is a vector $\mathbf{e}(c_{j}^{(i)})\in\{0,1\}^{K_{j}}$, where the $k$-th entry is defined as:
$\mathbf{e}(c_{j}^{(i)})_{k}=\begin{cases}1&\text{if }k=c_{j}^{(i)},\\ 0&\text{otherwise},\end{cases}$ (58)
for $k\in\{1,2,\ldots,K_{j}\}$. The one-hot encoded representation of $\mathbf{c}^{(i)}$ is then:
$\textsc{OneHot}(\mathbf{c}^{(i)})=[\mathbf{e}(c_{1}^{(i)}),\mathbf{e}(c_{2}^{(i)}),\ldots,\mathbf{e}(c_{D_{\text{cat}}}^{(i)})].$ (59)
The resulting vector has a total length of $\sum_{j=1}^{D_{\text{cat}}}K_{j}$, which corresponds to the sum of the unique categories across all categorical features. A softmax or a logarithm is then applied to the one-hot representation to yield a continuous probability distribution.
Learned Embeddings. Learned Embeddings (Learned) [34] encode the categorical component $\mathbf{c}^{(i)}$ using a representation trained directly within the model. Specifically, each categorical feature $c_{j}^{(i)}$ (the $j$-th element of $\mathbf{c}^{(i)}$) is assigned a trainable embedding vector of fixed dimensionality $d_{\text{EMB}}$. Hence, for each $j\in\{1,\dots,D_{\text{cat}}\}$, we have an embedding matrix
$E_{j}\;\in\;\mathbb{R}^{K_{j}\times d_{\text{EMB}}}.$
The embedding lookup operation retrieves the embedding vector for each $c_{j}^{(i)}$, and these vectors are then concatenated to form the full embedding for the categorical features:
$\textsc{Learned}(\mathbf{c}^{(i)})\;=\;\text{concat}\bigl{(}E_{1}[c_{1}^{(i)}],\;E_{2}[c_{2}^{(i)}],\;\ldots,\;E_{D_{\text{cat}}}[c_{D_{\text{cat}}}^{(i)}]\bigr{)}\;\in\;\mathbb{R}^{D_{\text{cat}}\cdot d_{\text{EMB}}}.$ (60)
To decode the learned embeddings back into categorical values, a nearest-neighbor approach is applied. For each embedding segment corresponding to a categorical feature, the pairwise distance between the embedding and the learned weights is computed, and the category with the minimum distance is selected:
$\hat{c}_{j}^{(i)}\;=\;\underset{k\in\{1,\dots,K_{j}\}}{\arg\min}\left\|\tilde{\mathbf{c}}_{j}^{(i)}-E_{j}[k]\right\|,$ (61)
where $\tilde{\mathbf{c}}_{j}^{(i)}\in\mathbb{R}^{d_{\text{EMB}}}$ is the segment of the concatenated embedding corresponding to the $j$-th categorical feature, and $E_{j}\in\mathbb{R}^{K_{j}\times d_{\text{EMB}}}$ is the embedding matrix for that feature.
Analog Bits. Analog Bits (i2b) [10] encode categorical features using a binary-based continuous representation. The encoding process involves two steps. We convert the categorical value to a real-valued binary representation where each category can be expressed using $\lceil\log_{2}(K)\rceil$ binary bits based on the number of categories:
[…]
followed by a shift and scale formula:
[…]
Thus, training and sampling of continuous-feature generative models (e.g., diffusion models) become computationally tractable. To decode, thresholding and rounding are applied to the generated continuous bits from the model to convert them back into binary form, which can be decoded trivially back into the original categorical values.
Dictionary. Dictionary encoding (dic) represents categorical features using a continuous look-up embedding function. The encoding assigns equally spaced real-valued representations within a tunable specified range, $[-1,1]$, to balance sparsity and separability. For categorical features with more categories, a wider range may be necessary to ensure proper distinction between values. The encoding process is defined as:
$\textsc{dic}(c_{j}^{(i)},K_{j}^{(i)})=-1+\frac{2c_{j}^{(i)}}{K_{j}^{(i)}-1},\quad\textsc{dic}(c_{j}^{(i)},K_{j}^{(i)})\in[-1,1]$ (64)
To decode, the nearest embedding is determined by comparing the encoded continuous value with all $K_{j}^{(i)}$ possible embeddings, selecting the category with the smallest Euclidean distance.
B.2 Architecture
Ordering of Categorical Feature. Our CatConverter representation induces an order on the categorical features. We assess how the order influences the results. By default, we assign a lexicographic ordering to the categorical features for simplicity purposes. We conducted an experiment to display the performance of lexicographic ordering vs. random ordering on the Adult and Beijing datasets. These datasets are chosen since they contain variables with a natural ordering such as “Education” (ordinal) and “Day” (periodic). As observed by the AUC disparity in Table 7, the order is an important factor in our CatConverter representation. It also implicitly highlights that CatConverter’s geometry aids in preserving the inherent semantics of ordinal and cyclical categorical features.
Table 7: AUC (classification) and RMSE (regression) scores of Machine Learning Efficiency. Higher scores indicate better performance.
AUC $\uparrow$ RMSE $\downarrow$
Methods Adult Beijing
TabRep-DDPM (Random) 0.776$\pm$.002 1.050$\pm$.006
TabRep-DDPM (Lexicographic) 0.913$\pm$.002 0.508$\pm$.006
TabRep-Flow (Random) 0.807$\pm$.003 1.041$\pm$.005
TabRep-Flow (Lexicographic) 0.912$\pm$.002 0.536$\pm$.006
Out-of-index (OOI). Out-of-index (OOI) can potentially occur due to the generative nature of DDPM and FM. For CatConverter, OOI values are cast to the value of the 00-th index. Although casting ensures that all generated categorical values fall within the valid range, it introduces a bias. We conducted additional experiments highlighting the casting rate that occurs when using our method. As observed in Table 8, casting rate is relatively low for most datasets, ranging between 5% to 20% apart from the Stroke dataset at around 30%. Nonetheless, we still achieve exceptional results across all datasets and benchmarks validating the effectiveness of our method.
Table 8: Out-of-Index Casting Rate
Methods Adult Default Shoppers Stroke Diabetes Beijing News
TabRep-DDPM 0.100$\pm$.001 0.094$\pm$.001 0.112$\pm$.005 0.272$\pm$.003 0.183$\pm$.002 0.072$\pm$.002 0.092$\pm$.003
TabRep-Flow 0.108$\pm$.001 0.069$\pm$.001 0.138$\pm$.001 0.348$\pm$.002 0.149$\pm$.001 0.054$\pm$.001 0.093$\pm$.001
Flow Matching/DDPM Denoising Network. The input layer projects the batch of tabular data input samples $\mathbf{z}_{t}$, each with dimension $d_{in}$, to the dimensionality $d_{t}$ of our time step embeddings $t_{emb}$ through a fully connected layer. This is so that we may leverage temporal information, which is appended to the result of the projection in the form of sinusoidal time step embeddings.
$h_{in}=\texttt{FC}_{d_{t}}(\mathbf{z}_{t})+t_{emb}$ (65)
Subsequently, the output is passed through hidden layers $h_{1}$, $h_{2}$, $h_{3}$, and $h_{4}$ which are fully connected networks used to learn the denoising direction or vector field. The output dimension of each layer is chosen as $d_{t}$, $2d_{t}$, $2d_{t}$, and $d_{t}$ respectively. On top of the FC networks, each layer also consists of an activation function followed by dropout, as seen in the formulas below. This formulation is repeated for each hidden layer, at the end of which we obtain $h_{out}$. The exact activations, dropout, and other hyperparameters chosen are shown in Table 9.
$h_{1}=\texttt{Dropout}(\texttt{Activation}(\texttt{FC}(h_{in})))$ (66)
At last, the output layer transforms $h_{out}$, of dimension $t_{emb}$ back to dimension $d_{in}$ through a fully connected network, which now represents the score function $\nabla_{\mathbf{z}_{t}}\log p_{\theta}(\mathbf{z}_{t})$.
$\nabla_{\mathbf{z}_{t}}\log p_{\theta}(\mathbf{z}_{t})=\texttt{FC}_{d_{in}}(h_{out})$ (67)
B.3 Hyperparameters
The hyperparameters selected for our model are shown in Table 9. The remaining hyperparameters of the baselines are tuned accordingly.
Table 9: TabRep Hyperparameters.
General Flow Matching/DDPM Denoising MLP
Hyperparameter Value Hyperparameter Value
Training Iterations $100,000$ Timestep embedding dimension $d_{t}$ $1024$
Flow Matching/DDPM Sampling Steps $50/1000$ Activation ReLU
Learning Rate $1\mathrm{e}{-4}$ Dropout $0.0$
[…]
Batch Size $4096$
Optimizer Adam
Appendix C Experiments
C.1 Datasets
Experiments were conducted with a total of 7 tabular datasets from the UCI Machine Learning Repository [13] with a (CC-BY 4.0) license. Classification tasks were performed on the Adult, Default, Shoppers, Stroke, and Diabetes datasets, while regression tasks were performed on the Beijing and News datasets. Each dataset was split into training, validation, and testing sets with a ratio of 8:1:1, except for the Adult dataset, whose official testing set was used and the remainder split into training and validation sets with an 8:1 ratio, and the Diabetes dataset, which was split into a ratio of 6:2:2. The resulting statistics of each dataset are shown in Table 10.
Table 10: Dataset Statistics. “# Num” and “# Cat” refer to the number of numerical and categorical columns.
Dataset # Samples # Num # Cat # Max Cat # Train # Validation # Test Task Type
Adult $48,842$ $6$ $9$ $42$ $28,943$ $3,618$ $16,281$ Binary Classification
Default $30,000$ $14$ $11$ $11$ $24,000$ $3,000$ $3,000$ Binary Classification
Shoppers $12,330$ $10$ $8$ $20$ $9,864$ $1,233$ $1,233$ Binary Classification
Beijing $41,757$ $7$ $5$ $31$ $33,405$ $4,175$ $4,175$ Regression
News $39,644$ $46$ $2$ $7$ $31,714$ $3,965$ $3,965$ Regression
Stroke $4,909$ $3$ $8$ $5$ $3,927$ $490$ $490$ Binary Classification
Diabetes $99,473$ $8$ $21$ $10$ $59,683$ $19,895$ $19,895$ Multiclass Classification
C.2 Baselines
TabRep’s performance is evaluated in comparison to previous works in diffusion-based mixed-type tabular data generation. This includes STaSy [24], CoDi [30], TabDDPM [26], TabSYN [58], and TabDiff [47]. The underlying architectures and implementation details of these models are presented in Table 11. Note that different benchmarks were used for CDTD [36] thus, we decided to not include it in our baselines.
Table 11: Comparison of tabular data synthesis baselines.
Method Model1 Type2 Categorical
Encoding
Numerical
Encoding
Additional Techniques
STaSy Score-based Diffusion U One-Hot
Encoding
Min-max scaler Self-paced learning and fine-tuning.
CoDi DDPM/ Multinomial
Diffusion
S One-Hot
Encoding
Min-max scaler Model Inter-conditioning and Contrastive learning to learn dependencies between categorical and numerical data.
TabDDPM DDPM/ Multinomial
Diffusion
S One-Hot
Encoding
Quantile
Transformer
Concatenation of numerical and categorical features.
TabSYN VAE + EDM U VAE-Learned Quantile
Transformer
Feature Tokenizer and Transformer encoder to learn cross-feature relationships with adaptive loss weighing to increase reconstruction performance.
TabDiff EDM/Masked Diffusion S One-Hot
Encoding
Quantile
Transformer
Joint continuous-time diffusion process of numerical and categorical variables under learnable noise schedules, with a stochastic sampler to correct sampling errors.
TabRep-DDPM DDPM U Cat-
Converter
Quantile
Transformer
Plug-and-play for Diffusion Models.
TabRep-Flow Flow
Matching
U Cat-
Converter
Quantile
Transformer
Plug-and-play for Diffusion Models.
  • 1
    The “Model” Column indicates the underlying architecture used for the model. Options include Denoising Diffusion Probabilistic Models or DDPMs [19], Multinomial Diffusion [21], EDM, as introduced in [23].
  • 2
    The “Type” column indicates the data integration approach used in the model. “U” denotes a unified data space where numerical and categorical data are combined after initial processing and fed collectively into the model. “S” represents a separated data space, where numerical and categorical data are processed and fed into distinct models.
C.3 Benchmarks
We evaluate the generative performance on a broad suite of benchmarks. We analyze the capabilities in downstream tasks such as machine learning efficiency (MLE), where we determine the AUC score for classification tasks and RMSE for regression tasks of XGBoost [9] on the generated synthetic datasets. Next, we conduct experiments on low-order statistics where we perform column-wise density estimation (CDE) and pair-wise column correlation (PCC). Lastly, we examine the models’ quality on high-order metrics such as $\alpha$-precision and $\beta$-recall scores [1]. We add three additional benchmarks including a detection test metric, Classifier Two Sample Tests (C2ST) [45], and privacy preservation metrics: the precision and recall of a membership inference attack (MIA) [48]. In this section, we expand on the concrete formulations behind our benchmarks including machine learning efficiency, low-order statistics, and high-order metrics. We also provide an overview on the detection and privacy metrics used in our experiments.
Machine Learning Efficiency. To evaluate the quality of our generated synthetic data, we use the data to train a classification/regression model, using XGBoost [9]. This model is applied to the real test set. AUC (Area Under Curve) is used to evaluate the efficiency of our model in binary classification tasks. It measures the area under the Receiver Operating Characteristic (or ROC) curve, which plots the True Positive Rate against the False Positive Rate. AUC may take values in the range [0,1]. A higher AUC value suggests that our model achieves a better performance in binary classification tasks and vice versa.
$\text{AUC}=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR})$ (68)
RMSE (Root Mean Square Error) is used to evaluate the efficiency of our model in regression tasks. It measures the average magnitude of the deviations between predicted values ($\hat{y}_{i}$) and actual values ($y_{i}$). A smaller RMSE model indicates a better fit of the model to the data.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (69)
Low-Order Statistics. Column-wise Density Estimation between numerical features is achieved with the Kolmogorov-Smirnov Test (KST). The Kolmogorov-Smirnov statistic is used to evaluate how much two underlying one-dimensional probability distributions differ, and is characterized by the below equation:
${\rm KST}=\sup\limits_{x}|F_{1}(x)-F_{2}(x)|,$ (70)
where $F_{n}(x)$, the empirical distribution function of sample n is calculated by
${\rm F_{n}(x)}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{(-\infty,x]}(X_{i})$ (71)
Column-wise Density Estimation between two categorical features is determined by calculating the Total Variation Distance (TVD). This statistic captures the largest possible difference in the probability of any event under two different probability distributions. It is expressed as
${\rm TVD}=\frac{1}{2}\sum\limits_{x\in X}|P_{1}(x)-P_{2}(x)|,$ (72)
where $P_{1}(x)$ and $P_{2}(x)$ are the probabilities (PMF) assigned to data point x by the two sample distributions respectively.
Pair-wise Column Correlation between two numerical features is computed using the Pearson Correlation Coefficient (PCC). It assigns a numerical value to represent the linear relationship between two columns, ranging from -1 (perfect negative linear correlation) to +1 (perfect positive linear correlation), with 0 indicating no linear correlation. It is computed as:
${\rho(x,y)}=\frac{{\rm cov}(x,y)}{\sigma_{x}\sigma_{y}},$ (73)
To compare the Pearson Coefficients of our real and synthetic datasets, we quantify the dissimilarity in pair-wise column correlation between two samples
Pearson ${\text{Pearson Score}}=\frac{1}{2}\mathbb{E}_{x,y}|\rho^{1}(x,y)-\rho^{2}(x,y)|$ (74)
Pair-wise Column Correlation between two categorical features in a sample is characterized by a Contingency Table. This table is constructed by tabulating the frequencies at which specific combinations of the levels of two categorical variables work and recording them in a matrix format.
To quantify the dissimilarity of contingency matrices between two different samples, we use the Contingency Score.
Contingency ${\text{Contingency Score}}=\frac{1}{2}\sum\limits_{\alpha\in A}\sum\limits_{\beta\in B}|P_{1,(\alpha,\beta)}-P_{2,(\alpha,\beta)}|,$ (75)
where $\alpha$ and $\beta$ describe possible categorical values that can be taken in features $A$ and $B$. $P_{1,(\alpha,\beta)}$ and $P_{2,(\alpha,\beta)}$ refer to the contingency tables representing the features $\alpha$ and $\beta$ in our two samples, which in this case corresponds to the real and synthetic datasets.
In order to obtain the column-wise density estimation and pair-wise correlation between a categorical and a numerical feature, we bin the numerical data into discrete categories before applying TVD and Contingency score respectively to obtain our low-order statistics.
We utilize the implementation of these experiments as provided by the SDMetrics library [^1] https://github.com/sdv-dev/SDMetrics.
High-Order Statistics. We utilize the implementations of High-Order Statistics as provided by the synthcity [^2] https://github.com/vanderschaarlab/synthcity library. $\alpha$-precision measures the overall fidelity of the generated data and is an extension of the classical machine learning quality metric of "precision". This formulation is based on the assumption that $\alpha$ fraction of our real samples are characteristic of the original data distribution and the rest are outliers. $\alpha$-precision therefore quantifies the percentage of generated synthetic samples that match $\alpha$ fraction of real samples. $\beta$-recall characterizes the diversity of our synthetic data and is similarly based on the quality metric of "recall". $\beta$-recall shares a similar assumption as $\alpha$-precision, except that we now assume that $\beta$ fraction of our synthetic samples are characteristic of the distribution. Therefore, this measure obtains the fraction of the original data distribution represented by the $\beta$ fraction of our generated samples [1].
Detection Metric: Classifier Two-Sample Test (C2ST). The Classifier Two-Sample Test, a detection metric, assesses the ability to distinguish real data from synthetic data. This is done through a machine learning model that attempts to label whether a data point is synthetic or real. The score ranges from 0 to 1 where a score closer to 1 is superior, indicating that the machine learning model cannot concretely identify whether the data point is real or generated. We select logistic regression as our machine learning model, using the implementation provided by SDMetric [45].
Privacy Metric: Membership Inference Attacks (MIA). Membership inference attacks evaluate the vulnerability of machine learning models to privacy leakage by determining whether a given instance was included in the training dataset [48]. The attacker often constructs a shadow model to mimic the target model’s behavior and trains a binary classifier to distinguish membership status based on observed patterns. We replaced DCR with Membership Inference Attacks since existing privacy ML literature [15, 54] conducted research highlighting the “Inadequacy of Similarity-based Privacy Metrics” such as DCR.
These attacks are evaluated using precision, the fraction of inferred members that are true members, and recall, the fraction of true members correctly identified. When records are equally balanced between members and non-members, the ideal precision and recall are $0.5$, indicating that the attack is no better than random guessing. Higher values suggest privacy leakage and reveal vulnerabilities in the model. Implementation of this metric is provided by SynthEval [28].
Appendix D Further Experimental Results
We perform experiments using several diffusion-based tabular generative model baselines, including STaSy [24], CoDi [30], TabDDPM [26], TabSYN [58]. We include TabDiff’s reported results as the authors have not released code [47].
We also incorporate several ablation experiments. We introduce TabFlow, an adaptation of TabDDPM [26] that models numerical and categorical tabular data with continuous flow matching [31, 33] and discrete flow matching [51], to examine the effect of unifying the data space. Experiments are also performed on various categorical representations on both diffusion and flow models. This includes one-hot, 1D-Learned Embedding, 2D-Learned Embedding, and Analog Bits [10], to demonstrate TabRep’s effectiveness. We show that our proposed TabRep framework achieves superior performance on the vast majority of metrics in Section D.1.
We evaluate AUC (classification), RMSE (regression), Column-Wise Density Estimation (CDE), Pair-Wise Column Correlation (PCC), $\alpha$-Precision, $\beta$-Recall scores, Classifier-Two Sample Test scores (C2ST), and Membership Inference Attacks Precision (MIA P.) and Recall (MIA R.) scores for our 7 datasets. $\uparrow$ indicates that the higher the score, the better the performance; $\downarrow$ indicates that the lower the score, the better the performance; $\updownarrow$ indicates that an optimal score should be as close to $50\%$ as possible.
The metrics and error bars shown in the tables in this section are derived from the mean and standard deviation of the experiments performed on $20$ sampling iterations on the best-validated model.
D.1 Additional Baseline Results
Adult
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.906$\pm$.001 92.26$\pm$0.04 89.15$\pm$0.10 77.05$\pm$0.29 33.54$\pm$0.36 55.37 49.52$\pm$0.50 24.51$\pm$0.44
CoDi 0.871$\pm$0.006 74.28$\pm$0.08 77.38$\pm$0.21 74.45$\pm$0.35 8.74$\pm$0.16 15.80 55.00$\pm$7.26 0.05$\pm$0.01
TabDDPM 0.910$\pm$0.001 98.37$\pm$0.08 96.69$\pm$0.09 90.99$\pm$0.35 62.19$\pm$0.48 97.55 51.30$\pm$0.07 56.90$\pm$0.29
TabSyn 0.906$\pm$0.001 98.89$\pm$0.03 97.56$\pm$0.06 98.97$\pm$0.26 47.68$\pm$0.27 95.49 50.91$\pm$0.16 42.91$\pm$0.31
TabDiff 0.912$\pm$0.002 99.37$\pm$0.05 98.51$\pm$0.16 99.02$\pm$0.20 51.64$\pm$0.20 99.50 51.03$\pm$0.12 52.00$\pm$0.35
TabRep-DDPM 0.913$\pm$0.002 99.39$\pm$0.04 98.63$\pm$0.04 99.11$\pm$0.25 52.04$\pm$0.12 99.50 50.44$\pm$0.83 52.78$\pm$0.21
TabRep-Flow 0.912$\pm$0.002 98.63$\pm$0.02 97.55$\pm$0.23 98.21$\pm$0.34 49.91$\pm$0.28 95.48 50.65$\pm$0.20 50.51$\pm$0.21
Default
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.752$\pm$0.006 89.41$\pm$0.03 92.64$\pm$0.03 94.83$\pm$0.15 40.23$\pm$0.22 62.82 51.84$\pm$0.69 30.37$\pm$0.99
CoDi 0.525$\pm$0.006 81.07$\pm$0.08 86.25$\pm$0.73 81.21$\pm$0.11 19.75$\pm$0.30 42.28 46.66$\pm$2.54 3.41$\pm$0.53
TabDDPM 0.761$\pm$0.004 98.20$\pm$0.05 97.16$\pm$0.19 96.78$\pm$0.30 53.73$\pm$0.28 97.12 51.34$\pm$0.49 44.96$\pm$0.59
TabSyn 0.755$\pm$0.005 98.61$\pm$0.08 98.33$\pm$0.67 98.49$\pm$0.20 46.06$\pm$0.37 95.83 50.80$\pm$0.56 43.71$\pm$0.89
TabDiff 0.763$\pm$0.005 98.76$\pm$0.07 97.45$\pm$0.75 98.49$\pm$0.28 51.09$\pm$0.25 97.74 51.15$\pm$0.62 46.67$\pm$0.55
TabRep-DDPM 0.764$\pm$0.005 98.97$\pm$0.19 96.74$\pm$0.62 98.66$\pm$0.24 48.22$\pm$0.48 98.90 50.07$\pm$0.41 48.96$\pm$0.41
TabRep-Flow 0.782$\pm$0.005 97.45$\pm$0.06 92.86$\pm$1.75 96.50$\pm$0.44 49.99$\pm$0.23 89.36 50.04$\pm$0.95 47.07$\pm$0.40
Shoppers
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.914$\pm$0.005 82.53$\pm$0.19 81.40$\pm$0.27 68.18$\pm$0.29 26.24$\pm$0.60 25.82 51.23$\pm$2.38 17.54$\pm$0.19
CoDi 0.865$\pm$0.006 67.27$\pm$0.03 80.52$\pm$0.12 90.52$\pm$0.37 19.22$\pm$0.27 19.04 59.67$\pm$11.74 1.36$\pm$0.45
TabDDPM 0.915$\pm$0.004 97.58$\pm$0.18 96.72$\pm$0.22 90.85$\pm$0.60 72.46$\pm$0.46 83.49 51.24$\pm$1.48 46.08$\pm$1.57
TabSyn 0.918$\pm$0.004 96.00$\pm$0.12 95.18$\pm$0.11 96.28$\pm$0.24 45.79$\pm$0.31 83.77 51.00$\pm$1.08 42.14$\pm$0.76
TabDiff 0.919$\pm$0.005 98.72$\pm$0.09 98.26$\pm$0.08 99.11$\pm$0.34 49.75$\pm$0.64 98.43 51.11$\pm$1.11 46.86$\pm$1.20
TabRep-DDPM 0.926$\pm$0.005 98.97$\pm$0.10 97.62$\pm$0.02 96.14$\pm$0.19 53.68$\pm$0.73 96.37 49.86$\pm$0.98 48.16$\pm$0.90
TabRep-Flow 0.919$\pm$0.005 97.74$\pm$0.03 97.08$\pm$0.07 95.85$\pm$0.46 55.92$\pm$0.37 94.20 51.38$\pm$1.66 49.19$\pm$0.86
Stroke
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.833$\pm$0.03 89.36$\pm$0.13 84.99$\pm$0.09 91.49$\pm$0.33 39.92$\pm$0.76 40.64 54.22$\pm$1.14 34.63$\pm$1.39
CoDi 0.798$\pm$0.032 87.42$\pm$0.17 80.65$\pm$1.81 86.46$\pm$0.53 28.59$\pm$0.47 24.47 54.54$\pm$1.60 27.32$\pm$3.50
TabDDPM 0.808$\pm$0.033 99.10$\pm$0.05 97.09$\pm$1.17 98.05$\pm$0.14 71.42$\pm$0.30 100.00 53.45$\pm$2.89 55.12$\pm$0.81
TabSyn 0.845$\pm$0.035 96.79$\pm$0.08 95.18$\pm$0.22 95.49$\pm$0.41 48.85$\pm$0.26 89.93 51.11$\pm$1.54 47.97$\pm$1.27
TabDiff 0.848$\pm$0.021 99.09$\pm$0.12 97.91$\pm$0.22 98.95$\pm$0.53 49.91$\pm$0.86 99.87 52.20$\pm$1.61 47.15$\pm$1.30
TabRep-DDPM 0.869$\pm$0.027 99.14$\pm$0.20 97.11$\pm$0.60 98.32$\pm$0.82 57.17$\pm$0.77 100.00 51.74$\pm$1.85 50.89$\pm$0.93
TabRep-Flow 0.854$\pm$0.028 98.42$\pm$0.31 97.37$\pm$2.12 96.40$\pm$0.71 63.91$\pm$0.87 95.96 50.66$\pm$1.77 49.43$\pm$1.16
Diabetes
Methods F1 $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.374$\pm$0.003 95.25$\pm$0.03 93.41$\pm$0.08 90.00$\pm$0.17 39.62$\pm$0.23 54.71 49.47$\pm$0.22 29.75$\pm$0.16
CoDi 0.288$\pm$0.009 76.42$\pm$0.02 78.07$\pm$0.18 38.96$\pm$0.16 6.39$\pm$0.16 3.95 0.00$\pm$0.00 0.00$\pm$0.00
TabDDPM 0.376$\pm$0.003 99.26$\pm$0.01 98.71$\pm$0.01 95.36$\pm$0.32 52.65$\pm$0.03 92.18 50.40$\pm$0.65 52.06$\pm$0.11
TabSyn 0.361$\pm$0.001 99.04$\pm$0.01 98.32$\pm$0.03 98.08$\pm$0.14 45.08$\pm$0.04 93.38 50.10$\pm$0.34 44.42$\pm$0.28
TabDiff 0.353$\pm$0.006 98.72$\pm$0.02 97.80$\pm$0.03 96.84$\pm$0.14 36.96$\pm$0.18 91.60 49.27$\pm$0.41 31.01$\pm$0.22
TabRep-DDPM 0.373$\pm$0.003 99.36$\pm$0.02 98.75$\pm$0.03 97.19$\pm$0.22 45.75$\pm$0.49 92.65 50.07$\pm$0.37 51.43$\pm$0.13
TabRep-Flow 0.377$\pm$0.002 99.00$\pm$0.02 98.46$\pm$0.05 99.08$\pm$0.13 48.58$\pm$0.17 90.41 50.24$\pm$0.37 50.33$\pm$0.13
Beijing
Methods RMSE $\downarrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.656$\pm$0.014 93.14$\pm$0.07 90.63$\pm$0.11 96.41$\pm$0.10 51.35$\pm$0.16 77.80 48.98$\pm$0.78 34.06$\pm$0.33
CoDi 0.818$\pm$0.021 83.54$\pm$0.04 90.35$\pm$0.21 96.89$\pm$0.14 53.16$\pm$0.12 80.27 39.19$\pm$6.35 0.40$\pm$0.09
TabDDPM 0.592$\pm$0.012 99.09$\pm$0.02 96.71$\pm$0.18 97.74$\pm$0.06 73.13$\pm$0.26 95.13 50.43$\pm$0.58 48.35$\pm$0.79
TabSyn 0.555$\pm$0.013 98.34$\pm$0.01 96.85$\pm$0.24 98.08$\pm$0.27 55.68$\pm$0.16 92.92 51.31$\pm$0.42 46.53$\pm$0.52
TabDiff 0.555$\pm$0.013 98.97$\pm$0.05 97.41$\pm$0.15 98.06$\pm$0.24 59.63$\pm$0.23 97.81 50.39$\pm$0.46 48.20$\pm$0.65
TabRep-DDPM 0.508$\pm$0.006 99.11$\pm$0.03 96.97$\pm$0.20 98.98$\pm$0.16 64.08$\pm$0.18 98.16 51.02$\pm$0.42 49.50$\pm$0.52
TabRep-Flow 0.536$\pm$0.006 98.28$\pm$0.07 96.92$\pm$0.21 98.16$\pm$0.13 62.65$\pm$0.12 92.26 50.35$\pm$0.60 49.14$\pm$0.81
News
Methods RMSE $\downarrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
STaSy 0.871$\pm$0.002 90.50$\pm$0.10 96.59$\pm$0.02 97.95$\pm$0.14 38.68$\pm$0.30 51.72 49.68$\pm$0.70 23.49$\pm$0.67
CoDi 1.21$\pm$0.005 70.82$\pm$0.01 95.44$\pm$0.05 86.03$\pm$0.12 35.01$\pm$0.13 9.35 40.00$\pm$24.49 0.04$\pm$0.02
TabDDPM 3.46$\pm$1.25 94.79$\pm$0.03 89.52$\pm$0.10 90.94$\pm$0.31 40.82$\pm$0.42 0.02 50.72$\pm$0.97 9.88$\pm$0.52
TabSyn 0.866$\pm$0.021 98.22$\pm$0.04 98.53$\pm$0.02 95.83$\pm$0.33 43.97$\pm$0.27 95.85 49.86$\pm$0.75 34.42$\pm$0.90
TabDiff 0.866$\pm$0.021 97.65$\pm$0.03 98.72$\pm$0.04 97.36$\pm$0.17 42.10$\pm$0.32 93.08 52.46$\pm$0.75 16.03$\pm$0.66
TabRep-DDPM 0.836$\pm$0.001 98.46$\pm$0.01 99.09$\pm$0.05 95.35$\pm$0.11 48.49$\pm$0.12 96.70 49.87$\pm$0.99 40.10$\pm$0.55
TabRep-Flow 0.814$\pm$0.002 96.89$\pm$0.03 98.34$\pm$0.29 90.91$\pm$0.25 51.75$\pm$0.16 88.13 50.90$\pm$0.93 35.48$\pm$0.78
D.2 Additional Ablation Results
Adult
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.476$\pm$0.057 48.94$\pm$0.12 38.05$\pm$0.08 17.44$\pm$0.21 0.70$\pm$0.01 1.89 38.67$\pm$19.02 0.03$\pm$0.02
Learned1D-DDPM 0.611$\pm$0.008 53.44$\pm$3.39 27.97$\pm$3.71 6.64$\pm$0.05 0.00$\pm$0.01 0.00 10.00$\pm$10.00 0.00$\pm$0.00
Learned2D-DDPM 0.793$\pm$0.003 62.69$\pm$1.16 38.53$\pm$1.54 7.11$\pm$0.21 0.02$\pm$0.03 0.00 38.00$\pm$5.61 0.03$\pm$0.01
i2b-DDPM 0.911$\pm$0.001 99.10$\pm$0.07 97.55$\pm$0.26 98.21$\pm$0.18 47.44$\pm$0.08 98.01 50.56$\pm$0.15 50.68$\pm$0.87
dic-DDPM 0.912$\pm$0.002 98.95$\pm$0.03 97.65$\pm$0.12 97.99$\pm$0.54 51.09$\pm$0.24 93.61 51.34$\pm$0.20 51.62$\pm$0.33
OneHot-Flow 0.895$\pm$0.003 90.66$\pm$0.07 84.32$\pm$0.10 88.35$\pm$0.15 30.64$\pm$0.13 38.88 51.08$\pm$0.60 13.49$\pm$0.36
Learned1D-Flow 0.260$\pm$0.014 60.00$\pm$3.13 36.20$\pm$3.61 6.82$\pm$0.08 0.02$\pm$0.03 0.00 24.67$\pm$10.41 0.02$\pm$0.01
Learned2D-Flow 0.126$\pm$0.012 62.37$\pm$3.03 38.94$\pm$4.20 6.64$\pm$3.49 0.05$\pm$0.04 0.00 28.19$\pm$9.68 0.04$\pm$0.01
i2b-Flow 0.911$\pm$0.001 98.23$\pm$0.09 97.14$\pm$0.32 99.54$\pm$0.31 48.87$\pm$0.16 92.18 50.70$\pm$0.31 49.68$\pm$0.79
dic-Flow 0.910$\pm$0.002 98.10$\pm$0.03 96.63$\pm$0.03 99.64$\pm$0.10 50.29$\pm$0.12 90.70 50.55$\pm$0.16 42.03$\pm$0.42
TabFlow 0.908$\pm$0.002 96.21$\pm$0.05 93.59$\pm$0.05 86.76$\pm$0.28 53.15$\pm$0.14 77.48 50.99$\pm$0.38 43.90$\pm$0.18
TabRep-DDPM 0.913$\pm$0.002 99.39$\pm$0.04 98.63$\pm$0.04 99.11$\pm$0.25 52.04$\pm$0.12 99.50 50.44$\pm$0.83 52.78$\pm$0.21
TabRep-Flow 0.912$\pm$0.002 98.63$\pm$0.02 97.55$\pm$0.23 98.21$\pm$0.34 49.91$\pm$0.28 95.48 50.65$\pm$0.20 50.51$\pm$0.21
Default
Methods AUC $\downarrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.557$\pm$0.052 50.88$\pm$0.10 50.88$\pm$0.07 4.13$\pm$0.05 0.15$\pm$0.01 0.11 40.00$\pm$24.49 0.08$\pm$0.05
Learned1D-DDPM 0.575$\pm$0.012 72.53$\pm$1.92 51.95$\pm$2.30 10.96$\pm$2.03 0.00$\pm$0.01 0.13 0.00$\pm$0.00 0.00$\pm$0.00
Learned2D-DDPM 0.290$\pm$0.009 71.11$\pm$0.32 50.79$\pm$0.36 11.99$\pm$2.54 0.01$\pm$0.01 0.06 15.33$\pm$11.62 0.11$\pm$0.08
i2b-DDPM 0.762$\pm$0.003 98.76$\pm$0.06 98.36$\pm$0.12 98.20$\pm$0.17 47.46$\pm$0.37 98.34 50.95$\pm$0.45 45.33$\pm$0.77
dic-DDPM 0.763$\pm$0.007 98.33$\pm$0.11 92.76$\pm$0.43 97.20$\pm$0.35 49.62$\pm$0.37 96.88 51.60$\pm$0.70 48.13$\pm$0.75
OneHot-Flow 0.759$\pm$0.005 91.95$\pm$0.05 88.04$\pm$1.51 93.12$\pm$0.31 30.50$\pm$0.19 69.14 51.40$\pm$1.23 12.32$\pm$0.60
Learned1D-Flow 0.438$\pm$0.009 73.52$\pm$0.82 53.01$\pm$0.21 20.03$\pm$5.43 0.00$\pm$0.01 0.53 10.00$\pm$10.00 0.05$\pm$0.05
Learned2D-Flow 0.709$\pm$0.008 66.33$\pm$4.86 44.11$\pm$6.75 2.33$\pm$12.11 0.01$\pm$0.02 0.01 10.00$\pm$10.00 0.03$\pm$0.03
i2b-Flow 0.763$\pm$0.004 97.67$\pm$0.13 94.65$\pm$1.28 97.42$\pm$0.57 49.15$\pm$0.48 90.04 50.66$\pm$0.33 43.09$\pm$1.25
dic-Flow 0.759$\pm$0.007 97.27$\pm$0.03 92.26$\pm$1.77 95.97$\pm$0.27 51.29$\pm$0.18 90.58 51.36$\pm$0.62 45.95$\pm$0.76
TabFlow 0.742$\pm$0.008 97.38$\pm$0.03 95.01$\pm$1.44 96.92$\pm$0.12 53.12$\pm$0.29 85.89 50.05$\pm$0.53 44.11$\pm$0.65
TabRep-DDPM 0.764$\pm$0.005 98.97$\pm$0.19 96.74$\pm$0.62 98.66$\pm$0.24 48.22$\pm$0.48 98.90 50.07$\pm$0.41 48.96$\pm$0.41
TabRep-Flow 0.782$\pm$0.005 97.45$\pm$0.06 92.86$\pm$1.75 96.50$\pm$0.44 49.99$\pm$0.23 89.36 51.02$\pm$0.95 47.07$\pm$0.40
Shoppers
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.799$\pm$0.126 90.37$\pm$0.14 84.61$\pm$0.10 90.67$\pm$0.26 37.76$\pm$0.57 54.59 51.00$\pm$0.97 28.54$\pm$1.49
Learned1D-DDPM 0.876$\pm$0.028 78.94$\pm$4.13 62.67$\pm$6.40 21.94$\pm$7.75 1.17$\pm$0.81 1.36 59.09$\pm$18.85 0.84$\pm$0.26
Learned2D-DDPM 0.103$\pm$0.011 70.97$\pm$3.06 51.21$\pm$4.37 8.43$\pm$5.81 0.05$\pm$0.09 0.09 0.00$\pm$0.00 0.00$\pm$0.00
i2b-DDPM 0.919$\pm$0.005 98.67$\pm$0.05 98.00$\pm$0.03 97.69$\pm$0.63 50.86$\pm$0.11 96.28 51.77$\pm$0.98 46.80$\pm$1.91
dic-DDPM 0.910$\pm$0.005 97.83$\pm$0.15 96.19$\pm$0.09 95.33$\pm$0.80 56.14$\pm$0.85 91.09 50.68$\pm$0.52 46.02$\pm$1.20
OneHot-Flow 0.910$\pm$0.006 92.84$\pm$0.08 91.50$\pm$0.14 87.57$\pm$0.44 48.77$\pm$0.69 65.08 49.64$\pm$1.69 34.56$\pm$0.80
Learned1D-Flow 0.134$\pm$0.015 76.25$\pm$2.16 60.05$\pm$3.32 10.78$\pm$7.79 0.10$\pm$1.00 0.01 40.00$\pm$24.49 0.13$\pm$0.08
Learned2D-Flow 0.868$\pm$0.007 71.93$\pm$0.94 53.29$\pm$1.34 16.58$\pm$1.72 0.28$\pm$0.23 0.07 57.33$\pm$20.50 0.65$\pm$0.34
i2b-Flow 0.910$\pm$0.005 97.61$\pm$0.11 97.20$\pm$0.11 97.24$\pm$0.66 54.88$\pm$0.26 91.83 51.56$\pm$1.01 45.11$\pm$1.20
dic-Flow 0.903$\pm$0.006 96.89$\pm$0.14 95.78$\pm$0.24 95.84$\pm$0.38 52.39$\pm$0.26 88.74 50.86$\pm$0.71 52.53$\pm$0.44
TabFlow 0.914$\pm$0.002 95.03$\pm$0.04 92.87$\pm$0.04 77.55$\pm$0.19 61.94$\pm$0.53 73.74 51.62$\pm$0.82 42.59$\pm$0.74
TabRep-DDPM 0.926$\pm$0.005 98.97$\pm$0.10 97.62$\pm$0.02 96.14$\pm$0.19 53.68$\pm$0.73 96.37 49.86$\pm$0.98 48.16$\pm$0.90
TabRep-Flow 0.919$\pm$0.005 97.74$\pm$0.03 97.08$\pm$0.07 95.85$\pm$0.46 55.92$\pm$0.37 94.20 51.38$\pm$1.66 49.19$\pm$0.86
Stroke
Methods AUC $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.797$\pm$0.033 98.74$\pm$0.11 97.65$\pm$0.10 96.78$\pm$1.08 56.27$\pm$0.22 98.62 53.08$\pm$2.50 49.76$\pm$1.24
Learned1D-DDPM 0.743$\pm$0.032 60.20$\pm$11.13 35.86$\pm$15.52 6.97$\pm$42.82 0.26$\pm$3.51 0.01 15.00$\pm$0.20 0.33$\pm$10.00
Learned2D-DDPM 0.850$\pm$0.035 59.11$\pm$14.59 33.77$\pm$19.97 6.67$\pm$36.39 0.87$\pm$2.42 0.02 18.33$\pm$0.33 0.49$\pm$13.02
i2b-DDPM 0.852$\pm$0.029 99.02$\pm$0.18 95.12$\pm$1.47 98.14$\pm$0.16 64.11$\pm$0.75 99.77 50.72$\pm$2.05 49.11$\pm$1.59
dic-DDPM 0.824$\pm$0.021 98.98$\pm$0.09 98.00$\pm$2.13 98.25$\pm$0.62 65.00$\pm$0.88 99.39 52.97$\pm$1.19 52.03$\pm$1.76
OneHot-Flow 0.812$\pm$0.029 94.17$\pm$0.08 90.45$\pm$0.08 81.87$\pm$0.55 49.23$\pm$0.56 64.82 49.32$\pm$1.47 39.51$\pm$1.25
Learned1D-Flow 0.142$\pm$0.033 75.48$\pm$6.55 56.60$\pm$9.05 71.63$\pm$35.72 0.45$\pm$0.13 0.16 47.00$\pm$0.55 1.14$\pm$20.22
Learned2D-Flow 0.180$\pm$0.030 54.02$\pm$13.09 28.95$\pm$17.18 5.67$\pm$24.97 0.21$\pm$0.67 0.00 20.00$\pm$0.16 0.16$\pm$20.00
i2b-Flow 0.797$\pm$0.027 98.72$\pm$0.11 94.65$\pm$1.33 98.26$\pm$0.27 65.23$\pm$0.26 97.19 52.72$\pm$2.26 52.03$\pm$1.18
dic-Flow 0.807$\pm$0.019 98.50$\pm$0.23 92.71$\pm$2.35 97.76$\pm$0.64 65.77$\pm$1.95 97.33 51.45$\pm$1.29 48.78$\pm$1.65
TabFlow 0.868$\pm$0.035 97.72$\pm$0.01 96.00$\pm$0.03 89.21$\pm$0.82 68.30$\pm$0.60 89.19 51.80$\pm$1.67 47.32$\pm$1.95
TabRep-DDPM 0.869$\pm$0.027 99.14$\pm$0.20 97.11$\pm$0.60 98.32$\pm$0.82 57.17$\pm$0.77 100.00 51.74$\pm$1.85 50.89$\pm$0.93
TabRep-Flow 0.854$\pm$0.028 98.42$\pm$0.31 97.37$\pm$2.12 96.40$\pm$0.71 63.91$\pm$0.87 95.96 50.66$\pm$1.77 49.43$\pm$1.16
Diabetes
Methods F1 $\uparrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.363$\pm$0.008 69.22$\pm$0.04 50.14$\pm$0.06 9.93$\pm$0.16 2.34$\pm$0.11 1.74 45.56$\pm$0.03 0.36$\pm$2.52
Learned1D-DDPM 0.179$\pm$0.010 63.40$\pm$4.59 39.94$\pm$5.34 0.00$\pm$0.00 0.00$\pm$0.00 0.00 0.00$\pm$0.00 0.00$\pm$0.00
Learned2D-DDPM 0.205$\pm$0.007 64.30$\pm$1.31 41.49$\pm$1.74 0.01$\pm$0.01 0.00$\pm$0.00 0.00 8.00$\pm$0.00 0.00$\pm$8.00
i2b-DDPM 0.370$\pm$0.008 99.38$\pm$0.01 98.84$\pm$0.03 97.70$\pm$0.13 46.83$\pm$0.39 93.98 50.08$\pm$0.39 51.63$\pm$0.12
dic-DDPM 0.375$\pm$0.006 99.50$\pm$0.02 99.12$\pm$0.01 98.92$\pm$0.13 48.34$\pm$0.18 95.03 50.37$\pm$0.25 54.48$\pm$0.74
OneHot-Flow 0.372$\pm$0.005 96.65$\pm$0.03 94.61$\pm$1.99 97.43$\pm$0.06 41.64$\pm$0.16 55.44 49.49$\pm$0.21 22.78$\pm$0.26
Learned1D-Flow 0.184$\pm$0.007 53.27$\pm$7.48 28.31$\pm$8.51 0.00$\pm$0.00 0.00$\pm$0.00 0.00 0.00$\pm$0.00 0.00$\pm$0.00
Learned2D-Flow 0.177$\pm$0.008 68.39$\pm$9.23 46.71$\pm$10.95 0.00$\pm$0.00 0.00$\pm$0.00 0.00 0.00$\pm$0.00 0.00$\pm$0.00
i2b-Flow 0.372$\pm$0.003 98.92$\pm$0.03 98.34$\pm$0.05 98.45$\pm$0.14 48.97$\pm$0.29 89.28 49.86$\pm$0.26 48.25$\pm$0.13
dic-Flow 0.376$\pm$0.007 99.01$\pm$0.03 98.54$\pm$0.03 99.65$\pm$0.10 48.98$\pm$0.10 89.06 50.10$\pm$0.12 48.01$\pm$0.18
TabFlow 0.376$\pm$0.006 98.04$\pm$0.02 96.82$\pm$0.01 79.60$\pm$0.09 51.58$\pm$0.04 73.35 50.19$\pm$0.39 44.14$\pm$0.15
TabRep-DDPM 0.373$\pm$0.003 99.36$\pm$0.02 98.75$\pm$0.03 97.19$\pm$0.22 46.19$\pm$0.49 92.65 50.07$\pm$0.37 51.43$\pm$0.13
TabRep-Flow 0.377$\pm$0.002 99.00$\pm$0.02 98.46$\pm$0.05 99.08$\pm$0.13 48.58$\pm$0.17 90.41 50.24$\pm$0.37 50.33$\pm$0.13
Beijing
Methods RMSE $\downarrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 2.143$\pm$0.339 74.21$\pm$0.07 63.68$\pm$0.05 48.88$\pm$0.03 19.07$\pm$0.13 19.68 50.28$\pm$2.95 5.44$\pm$0.50
Learned1D-DDPM 0.921$\pm$0.006 79.49$\pm$3.50 62.23$\pm$5.61 26.94$\pm$27.91 8.49$\pm$5.95 13.51 45.45$\pm$2.82 1.97$\pm$0.13
Learned2D-DDPM 0.969$\pm$0.005 81.89$\pm$1.24 66.02$\pm$2.37 79.49$\pm$16.01 7.73$\pm$1.53 55.83 49.20$\pm$1.89 2.05$\pm$0.10
i2b-DDPM 0.542$\pm$0.008 98.66$\pm$0.04 96.95$\pm$0.21 97.92$\pm$0.15 59.27$\pm$0.14 94.28 50.64$\pm$0.47 48.43$\pm$0.89
dic-DDPM 0.547$\pm$0.007 98.83$\pm$0.04 97.21$\pm$0.16 98.97$\pm$0.30 61.90$\pm$0.12 98.83 51.32$\pm$0.35 50.61$\pm$0.66
OneHot-Flow 0.765$\pm$0.016 84.61$\pm$0.02 67.28$\pm$4.64 84.38$\pm$0.61 20.32$\pm$0.19 35.76 51.44$\pm$1.49 8.12$\pm$0.43
Learned1D-Flow 0.806$\pm$0.009 80.14$\pm$1.60 64.19$\pm$2.33 81.15$\pm$3.68 13.11$\pm$0.38 20.40 44.02$\pm$3.58 1.13$\pm$0.09
Learned2D-Flow 0.787$\pm$0.007 79.50$\pm$0.85 63.04$\pm$1.01 43.00$\pm$4.90 7.55$\pm$4.05 14.58 54.23$\pm$1.72 2.01$\pm$0.09
i2b-Flow 0.543$\pm$0.007 98.08$\pm$0.04 96.87$\pm$0.43 96.83$\pm$0.12 60.58$\pm$0.19 91.66 49.64$\pm$0.47 45.77$\pm$1.07
dic-Flow 0.561$\pm$0.013 98.09$\pm$0.03 96.37$\pm$0.08 97.04$\pm$0.22 60.78$\pm$0.10 93.96 50.86$\pm$0.71 52.53$\pm$0.44
TabFlow 0.574$\pm$0.01 96.44$\pm$0.06 93.71$\pm$0.07 94.81$\pm$0.42 59.47$\pm$0.28 87.23 50.54$\pm$0.35 42.20$\pm$0.61
TabRep-DDPM 0.508$\pm$0.006 99.11$\pm$0.03 96.97$\pm$0.20 98.98$\pm$0.16 64.08$\pm$0.18 98.16 51.02$\pm$0.42 49.50$\pm$0.52
TabRep-Flow 0.536$\pm$0.006 98.28$\pm$0.07 96.92$\pm$0.21 98.16$\pm$0.13 62.65$\pm$0.12 92.26 50.35$\pm$0.60 49.14$\pm$0.81
News
Methods RMSE $\downarrow$ CDE $\uparrow$ PCC $\uparrow$ $\alpha$uparrow $\beta$uparrow C2ST $\uparrow$ MIA P. $\updownarrow$ MIA R.$\updownarrow$
OneHot-DDPM 0.840$\pm$0.02 98.11$\pm$0.06 92.78$\pm$0.08 96.33$\pm$0.24 46.87$\pm$0.28 95.80 50.25$\pm$0.46 32.42$\pm$0.80
Learned1D-DDPM 0.858$\pm$0.01 96.45$\pm$0.14 95.00$\pm$0.32 90.48$\pm$9.11 19.54$\pm$4.84 21.58 51.33$\pm$0.96 17.60$\pm$1.02
Learned2D-DDPM 0.857$\pm$0.023 96.06$\pm$0.49 94.49$\pm$0.84 88.83$\pm$4.52 5.11$\pm$5.24 20.41 50.85$\pm$2.97 3.83$\pm$0.39
i2b-DDPM 0.844$\pm$0.013 98.41$\pm$0.02 98.57$\pm$0.18 95.02$\pm$0.10 48.22$\pm$0.18 96.07 50.88$\pm$0.57 39.82$\pm$0.93
dic-DDPM 0.866$\pm$0.019 98.22$\pm$0.06 98.50$\pm$0.36 97.08$\pm$0.25 47.88$\pm$0.10 95.75 49.54$\pm$0.57 39.80$\pm$0.35
OneHot-Flow 0.850$\pm$0.017 96.27$\pm$0.05 98.11$\pm$0.02 97.78$\pm$0.13 43.06$\pm$0.62 84.56 50.45$\pm$2.02 14.84$\pm$0.65
Learned1D-Flow 0.873$\pm$0.007 95.07$\pm$0.14 95.07$\pm$0.12 89.16$\pm$5.89 18.17$\pm$6.18 79.17 49.85$\pm$1.69 10.28$\pm$0.47
Learned2D-Flow 0.866$\pm$0.005 94.85$\pm$0.38 94.99$\pm$0.89 80.66$\pm$15.23 14.63$\pm$4.14 78.96 49.59$\pm$1.54 14.19$\pm$0.69
i2b-Flow 0.847$\pm$0.014 96.64$\pm$0.05 98.38$\pm$0.34 88.39$\pm$0.11 51.85$\pm$0.24 89.47 51.12$\pm$0.71 36.35$\pm$0.53
dic-Flow 0.853$\pm$0.014 96.58$\pm$0.04 97.49$\pm$0.34 92.28$\pm$0.25 50.79$\pm$0.29 88.30 50.43$\pm$0.63 37.56$\pm$0.77
TabFlow 0.850$\pm$0.017 96.51$\pm$0.08 97.93$\pm$0.02 92.68$\pm$0.31 50.03$\pm$0.26 87.33 51.01$\pm$0.24 28.00$\pm$0.89
TabRep-DDPM 0.836$\pm$0.001 98.46$\pm$0.01 99.09$\pm$0.05 95.35$\pm$0.11 48.49$\pm$0.12 96.70 49.87$\pm$0.99 40.10$\pm$0.55
TabRep-Flow 0.814$\pm$0.002 96.89$\pm$0.03 98.34$\pm$0.29 90.91$\pm$0.25 51.75$\pm$0.16 88.13 50.90$\pm$0.93 35.48$\pm$0.78
D.3 Additional Training and Sampling Duration Results
We conducted an additional Training and Sampling Duration experiment on the largest dataset among our dataset suite (Diabetes dataset) with $99,473$ samples, $8$ numerical features, and $21$ categorical features. As observed in Table 12, we save around $1700$ seconds compared to TabDDPM and around $4900$ seconds compared to TabSYN during training and sampling.
Table 12: Training and Sampling Duration in Seconds.
Methods Training Sampling Total
TabDDPM 3455 268 3723
TabSYN 6003 $+$ 882 18 6903
TabRep-DDPM 1980 114 2094
TabRep-Flow 2002 7 2009
D.4 Additional Results on High Cardinality and Imbalanced Toy Datasets
We curate synthetic toy datasets of high cardinality categorical variables and imbalanced datasets to reinforce our generalizability claims. Our high cardinality toy dataset is a regression task with two categorical features. The first categorical feature is of high cardinality, with $1000$ unique categories where each category is assigned a base effect drawn from a normal distribution. The other categorical feature may take $3$ values, each having fixed effects of $1.0$, $-1.0$, and $0.5$ respectively. The target label is computed by summing the base effect from the high-cardinality category, the fixed effect from the other categorical feature, and an independent numerical feature drawn from a standard normal distribution. Additional Gaussian noise is added to perturb the data.
Table 13: Performance on High Cardinality Setting.
RMSE $\downarrow$ CDE $\uparrow$ PWC $\uparrow$ C2ST $\uparrow$ $\alpha$-Precision $\uparrow$ $\beta$-Recall $\uparrow$
TabDDPM 0.8253$\pm$0.1419 42.93$\pm$0.17 11.20$\pm$0.09 0.41$\pm$0.02 0.46$\pm$0.01 0.02$\pm$0.01
TabSYN 0.4775$\pm$0.0129 94.23$\pm$0.56 78.37$\pm$2.00 100.00$\pm$0.00 99.13$\pm$0.31 35.35$\pm$0.47
TabRep-DDPM 0.4662$\pm$0.0071 80.88$\pm$0.15 59.96$\pm$1.08 25.45$\pm$0.10 99.31$\pm$0.21 16.47$\pm$0.32
TabRep-Flow 0.4812$\pm$0.0104 94.38$\pm$0.28 76.84$\pm$1.08 98.59$\pm$0.94 99.32$\pm$0.22 36.00$\pm$0.24
As shown in Table 13, it is worth noting that DDPM models including TabRep-DDPM and TabDDPM perform poorly in CDE, PWC, and C2ST tasks, yet are able to model RMSE and $\alpha$-precision well. This indicates that with high cardinality, TabRep-DDPM is less capable of learning conditional distributions across features. In contrast, our proposed TabRep-Flow performs on par with TabSYN, as flow-matching models’ smooth differentiable transformation allows them to capture subtle conditional variations, and TabSYN’s latent space allows for the learning of a simpler latent distribution.
The imbalanced toy dataset is a regression task with one binary categorical feature (distributed $95$% class A, $5$% class B). Each row also contains a numeric feature drawn from a standard normal distribution. The target is constructed by applying a category-specific linear function before addition of some Gaussian noise to generate variations in the data.
Table 14: Performance on Imbalanced Setting.
RMSE $\downarrow$ CDE $\uparrow$ PWC $\uparrow$ C2ST $\uparrow$ $\alpha$-Precision $\uparrow$ $\beta$-Recall $\uparrow$
TabDDPM 0.1694$\pm$0.0016 98.93$\pm$0.54 95.87$\pm$5.43 98.98$\pm$1.57 98.96$\pm$0.68 50.29$\pm$0.49
TabSYN 0.1708$\pm$0.0018 94.98$\pm$0.04 96.36$\pm$0.18 90.27$\pm$0.65 95.84$\pm$1.13 48.89$\pm$0.23
TabRep-DDPM 0.1688$\pm$0.0010 99.44$\pm$0.13 96.82$\pm$3.97 99.42$\pm$0.54 99.46$\pm$0.08 50.96$\pm$0.77
TabRep-Flow 0.1689$\pm$0.0025 98.52$\pm$0.12 90.49$\pm$5.36 99.06$\pm$0.47 96.69$\pm$0.28 51.30$\pm$0.31
In Table 14, we see that for imbalanced data, our proposed methods achieve results that are better compared to existing models like TabDDPM and TabSYN, showing that our methods are generalizable to cases where training data may be highly imbalanced.
D.5 TabSYN’s Latent Representation Dimension
By default, TabSYN has a latent dimensionality of 4. To address the concern regarding TabSYN’s dimensionality, we run TabSYN using the same dimensions (2D latent space) as our TabRep representation. As observed in Table 15, TabSYN with a 2D latent dimension performs much worse than TabSYN with a 4D latent dimension.
Table 15: AUC (classification) and RMSE (regression) scores of Machine Learning Efficiency. Higher scores indicate better performance.
AUC $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$
Methods Adult Default Shoppers Stroke Diabetes Beijing News
TabSYN 0.906$\pm$.001 0.755$\pm$.004 0.918$\pm$.004 0.845$\pm$.035 0.361$\pm$.001 0.586$\pm$.013 0.862$\pm$.021
TabSYN (2D Latent Space) 0.892$\pm$.002 0.752$\pm$.005 0.916$\pm$.002 0.811$\pm$.032 0.368$\pm$.002 0.720$\pm$.015 0.868$\pm$.003
TabRep-DDPM 0.913$\pm$.002 0.764$\pm$.005 0.926$\pm$.005 0.869$\pm$.027 0.373$\pm$.003 0.508$\pm$.006 0.836$\pm$.001
TabRep-Flow 0.912$\pm$.002 0.782$\pm$.005 0.919$\pm$.005 0.830$\pm$.028 0.377$\pm$.002 0.536$\pm$.006 0.814$\pm$.002
In terms of computational costs on the Adult dataset, Table 16 highlights that TabSYN in a 2D Latent Space saves close to 500 seconds while compromising on accuracy when compared to vanilla TabSYN (4D Latent Space). However, it still consumes around 1000 seconds extra when compared to TabRep methods.
Table 16: Training and Sampling Duration in Seconds.
Methods Training Sampling Total
TabSYN 2374 $+$ 1085 11 3470
TabSYN (2D Latent Space) 2333 $+$ 670 6 3009
TabRep-DDPM 2071 59 2130
TabRep-Flow 2028 3 2031